### v1

# Crossroads of Continents: Automated Artifact Extraction for Cultural Adaptation with Large Multimodal Models

Abstract
In this work, we present a comprehensive three-phase study to examine (1) the effectiveness of large multimodal models (LMMs) in recognizing cultural contexts; (2) the accuracy of their representations of diverse cultures; and (3) their ability to adapt content across cultural boundaries. We first introduce Dalle Street, a large-scale dataset generated by DALL-E 3 and validated by humans, containing $9,935$ images of $67$ countries and $10$ concept classes. We reveal disparities in cultural understanding at the sub-region level with both open-weight (LLaVA) and closed-source (GPT-4V) models on Dalle Street and other existing benchmarks. Next, we assess models’ deeper culture understanding by an artifact extraction task and identify over $18,000$ artifacts associated with different countries. Finally, we propose a highly composable pipeline, CultureAdapt, to adapt images from culture to culture. Our findings reveal a nuanced picture of the cultural competence of LMMs, highlighting the need to develop culture-aware systems. [^1] Dataset and code are available: https://github.com/iamshnoo/crossroads
Crossroads of Continents: Automated Artifact Extraction for Cultural Adaptation with Large Multimodal Models
Anjishnu Mukherjee  Ziwei ZhuAntonios Anastasopoulos Department of Computer Science, George Mason University {amukher6,zzhu20,antonis}@gmu.edu
1 Introduction
Figure 1: We introduce a large scale balanced dataset for measuring cultural awareness, a task for extracting implicit cultural artifacts from images and a highly composable pipeline for adapting images to different cultural contexts with fine-grained edits.
Culture is hard to define and has always been so. Kroeber (1952) explored how the word has evolved to gain different meanings in different contexts. Recent efforts in natural language processing research have seen growing interest in understanding how culture influences language models and human behavior, including language, art, and decision-making Hershcovich et al. (2022); Adilazuarda et al. (2024); Liu et al. (2024); Ge et al. (2024). As large multimodal models (LMMs; Liu et al., 2023a) intersect more with human life, the need for them to comprehend and respect cultural nuances is crucial. Research in this area explores model alignment with human values, measures of cultural awareness, and methods for cultural adaptation, where content that may be considered as accurately representing a culture or a country, maybe stereotypically, is modified to represent a different one.
The challenge of assessing the capability of understanding and leveraging cultural knowledge in LMMs, given their multimodal nature – processing both text and image data – is significant. Prior research has primarily investigated LMMs for cultural awareness [^2] We maintain that LMMs do not inherently possess human values but that their outputs may display cultural knowledge. by examining their performance on culture-related tasks such as region classification from images Basu et al. (2023); Yin et al. (2023); Pouget et al. (2024), image-caption matching Liu et al. (2021), and cultural image captioning Cao et al. (2024). However, these tasks may not fully determine whether LMMs are responding to cultural cues encoded within their training data or merely identifying superficial culture-associated features. Furthermore, there is a lack of in-depth analysis of how LMMs utilize cultural knowledge during these tasks and whether they exhibit consistent understanding across various cultural contexts.
To address these gaps, We propose a comprehensive three-phase study to deepen our understanding of cultural awareness in LMMs. First, we develop a large-scale dataset for assessing cultural awareness as measured by the ability of LMMs to recognize and differentiate between cultures, with countries as proxies, in a task setting similar to GeoGuessr Geoguessr (2024). Next, we delve deeper into cultural understanding by introducing an artifact extraction task designed to identify implicit cultural artifacts that LMMs use to distinguish between cultures. Finally, based on the insights gained from the artifact extraction task, we propose a cultural adaptation task combining multiple generative models in an end-to-end pipeline to adapt images from one cultural context to another. This pipeline enhances our understanding of LMMs’ adaptability and promises an effective data augmentation technique to improve cultural awareness in LMMs.
Our main contributions are as follows:
  • We introduce, Dalle Street, a collection of $9,935$ images generated by DALL-E 3, covering $67$ countries and $10$ cultural concept classes. Compared to datasets like Dollar Street Rojas et al. (2022), Dalle Street includes more images from underrepresented geographic regions. We validate the appropriateness of country representations with a human study and will release this dataset publicly.
  • We measure how well both open-weight and closed-source LMMs perform on Dalle Street based on the cultural knowledge encoded in them to understand disparities in performance at the geographic sub-region level for a diverse group of concepts and countries. We include a human study baseline to understand where current models may perform better than humans. We also compare LMM performance on two image datasets collected from different sources.
  • We introduce a task for cultural artifact extraction to identify implicit associations. We further develop a method to filter these effectively to discover interesting associations that frequently co-occur for each country.
  • We propose a highly composable end-to-end pipeline, CultureAdapt, using a chain of LMMs to identify a country for a given image, extract culturally relevant artifacts, ground them with bounding boxes, and use diffusion-based inpainting to adapt to a target culture. We show examples of this pipeline and include a CLIPScore-based metric to validate its performance.
2 Data
We use three different datasets covering a wide variety of cultural concepts, economic ranges, and different sources: (a) Dalle Street: synthetic, DALL-E 3 (OpenAI, 2024) generated; (b) Dollar Street: natural, collected photographs; and (c) MaRVL: scraped from the internet under native speaker guidance. We include count statistics for each of the datasets in Table C.2. Overall, our study uses roughly $20$k images across $3$ datasets, covering $19$ geographical sub-regions and $67$ countries from the United Nations geoscheme UN (2024). Each image is accompanied by a true label for country.
Dalle Street
We consider the same $10$ categories and $19$ geographical regions as in Dollar Street and include $67$ countries. [^3] The four extra countries corresponding to new additions to our annotator pool over the timeline of this project. Given these parameters, we generate high resolution images of size $1024$x1024102410241024 from DALL-E 3 OpenAI (2024) for 2 settings - vivid (for generating hyper-real and dramatic images) and natural (for more realistic images). To do so, we follow a templatic prompting approach (prompt shown in Appendix Figure 9) and then sample at least $10$ images for each combination of country and concept class. However, due to global content policy filters imposed on the API calls and limited budget, we end up with $9,935$ images (as opposed to the expected $13,400$). We will release this generated dataset under CC BY-SA 4.0.
We perform a qualitative study, where we ask annotators to determine the “appropriateness” of the generated image on a Likert scale from $1$ (Strongly Disagree) to $5$ (Strongly Agree), where we define appropriateness as the case where the image shown to the annotator is one of the (possibly stereotypical) representations for the given combination of country and category.
Dollar Street Rojas et al. (2022)
This is a dataset of photos of objects and scenes collected by professional and volunteer photographers which specifically focuses on poor and remote environments, containing approximately $200$ concept classes and covering $63$ countries. We filter it for images that do not contain multiple labels, classes which do not cover images for all regions, and classes with subjective naming. Then we refer to our group of annotators from diverse backgrounds to choose the top $10$ categories by the method of collaborative labeling Chang et al. (2017), where effectively we simplify our selection of object classes by choosing the ones which all annotators universally agree on as being a relevant dimension for testing cultural awareness. The final data subset that we end up with thus includes $63$ countries, $19$ geographical regions, $4,137$ images, $10$ concept classes ( car, cups, mugs and glasses, family snapshots, front door, home, kitchen, plate of food, social drink, wall decoration, and wardrobe). This dataset is available under CC BY-SA 4.0.
MaRVL Liu et al. (2021)
The primary task for this dataset involves validation of statements about image pairs for five languages: Indonesian, Mandarin Chinese, Swahili, Tamil, and Turkish. We use the images and assign country and region labels to create our task setting for cultural awareness. [^4] MaRVL language to region mapping in Appendix C.1 This dataset gives us $4,914$ images curated by native speakers of each of these languages, across $5$ distinct geographical regions and it is distributed under CC BY 4.0 license for research purposes.
3 Cultural Awareness (Task 1)
To measure cultural awareness of LMMs, we compare performance on two existing culturally diverse benchmarks and also on Dalle Street. Later on, we also compare this with the baseline from our human study to understand where LMMs stand at this task in comparison to humans. Overall, we find performance varies across sub-regions but both LLaVA and GPT-4V perform nearly equally well.
3.1 Methods
Given an input image, we prompt two LMMs, LLaVA-NeXT (Liu et al., 2023a), an open-weight model, and GPT-4V vision-preview (OpenAI et al., 2023), a closed-source model, in a zero-shot setting by asking an open-ended question without any possible answer choices: predict the geographical region represented in the image, as per the United Nations geoscheme. We use this geoscheme for three reasons. First, the refusal rate of models is a lot higher when asked about specific country labels. Second, the geoscheme information is already present on English Wikipedia, which is usually included in pre-training corpora for most LLMs. And third, sticking to a geoscheme allows us to parse open-ended generations in a structured format. A potential challenge with this approach is that models may not incorporate the most recent changes in the geoscheme, leading to finer grained errors which would be hard to trace (for example, the Wikipedia page for the geoscheme has been updated more than $20$ times since the cut-off date for the training data of vision-preview, including some changes to island groups like Channel Islands). We focus on countries which have traditionally been consistently classified without changes to geographic region to avoid these inconsistencies.
GPT-4V LLaVA
Dollar Street $36.28$ $\mathbf{36.83}$
Dalle Street $56.31$ $\mathbf{78.05}$
MaRVL $\mathbf{41.59}$ $19.14$
Table 1: Overall, LLaVA performs as good as or better than GPT-4V on two out of three datasets. Human accuracy on a subset of Dalle Street is $42.63$.
Metric: Accuracy
We process generated text to map it to one of the geographical sub-regions or to a policy violation case, and then compare with true labels by mapping country information to geographical regions, which gives us classification accuracy as a quantitative metric for measuring success. Specifically, we inspect the confusion matrix to better understand on which regions the models make mistakes.
Economic disparities
[…]
Human Study - GeoGuessr
We develop an annotation interface similar to Geoguessr (2024) and ask our pool of annotators to label images at either the country or the region level or at continent level. We ask for a maximum of $5$ labels and a minimum of $1$ per image, to adjust for different familiarity levels with multiple randomly sampled regions/cultures. Appendix C.3 includes detailed descriptions of annotator demographics, interface and instructions provided to annotators.
3.2 Results
Contrary to our expectations, LLaVA is not far off from GPT-4V in terms of cultural awareness, as measured in our experimental settings for this task.
Overall comparison
We show the aggregated results for our Cultural Awareness Task in Table 1. LLaVA performs as good as GPT-4V on Dollar Street and even outperforms it significantly on the Dalle Street images. This is a potential indication that LLaVA might be more prone to forming stereotypical associations between regions and concepts, because the Dalle Street images include such associations to distinguish different geographical regions. As we discuss later, it might be possible to trace this down to the training data.
However, we note that on the MaRVL data, LLaVA performs about as good as random guessing, specifically getting incorrect predictions for nearly all of the images corresponding to Swahili and Turkish. Our hypothesis to explain this is based on the fact that the other two datasets have similar distributions about concepts and countries, whereas MaRVL is much more diverse and covers fewer countries but many indigenous concepts, not all of which the model may have seen during training.
Figure 2: Confusion matrices for GPT-4V on the cultural awareness task for Dalle Street images. Accurate responses match the true subregion exactly. Special labels include Invalid (no match to region labels or incomplete) and ResponsibleAI (no region label due to policy violation). Takeaway: The model shows balanced performance, indicated by a strong leading diagonal. Notably, the model achieves $100$% accuracy for Western Asia images, covering Iran, Jordan, Lebanon, Oman, Palestine, and Turkey.
Subregion level analysis
Figure  2 (and Figure  C.4.15) indicates that both GPT-4V and LLaVA have strong leading diagonals, which is an indicator that they get many predictions correct. GPT-4V has many cases where it does not provide any answer due to violations of the content policy filter. Instead of trying to find workarounds for the filter, we show it as is, because this would represent the experience of any typical user of the system. Very interestingly, both GPT-4V and LLaVA gets all of the answers from Western Asia correct. In general, LLaVA makes a lot of mistakes for different regions, but defaults to South America as the incorrect answer for these cases, whereas for GPT-4V the same translates to the ResponsibleAI policy filter output as the incorrect answer. We include results for the other two datasets in the Appendix (Figure C.4.16 and  C.4.17 ), and find similar trends.
Figure 3: We normalize income data from Dollar Street into region specific quartiles and plot corresponding accuracies for GPT-4V. Takeaway: Lower income quartiles (blue, orange bars) are more accurate in Africa and Asia, but higher income quartiles (red, green) are more accurate in Americas; in Europe all quartiles have similar accuracies.
Economic disparity - GPT-4V
We utilize the monthly income information available as part of the Dollar Street data, and divide the images into groups based on normalized income quartiles for each broad geographical region - Africa, Asia, Americas, Europe and Oceania (see Figure 3). Lower income quartiles show better performance for Africa and Asia, whereas for America performance improves for higher income groups. This discrepancy reveals the latent biases of the model, which seem to associate Africa with poorer contexts and America with richer ones. For Europe, performance is nearly similar for all quartiles.
4 Automated Extraction of Implicit Cultural Artifacts (Task 2)
To identify the implicit associations that the models may be using for performing Task $1$, we propose a second task to extract specific cultural artifacts from the images, and discover interesting associations, which are usually stereotypical for the relevant countries.
Figure 4: We assign a score for each identified artifact based on how likely it is to co-occur for a given country. Across all such scores, we find the mean and standard deviation to define a range. Scores that lie outside this range would correspond to items that co-occur frequently for a given country, and would most likely represent stereotypical implicit associations.
4.1 Methods
We use the open vocabulary object detection Zareian et al. (2021) capabilities of GPT-4V vision-preview, formulating a detailed prompt (see Appendix C.2) to extract information about the target concept class in a Dalle Street image, including text descriptions, color, [^5] Note that this corresponds to appearance of different objects and persons in images, and does not imply race. and person count. GPT-4V’s strong instruction following capabilities allow the generation of a nearly perfectly validated JSON string, which we lightly post-process, and then provide that as input to GPT-4 turbo to get a summary of the objects for any given image. In our initial experiments, GPT-4V outperformed LLaVA significantly in terms of following complex instructions for object detection, so we only report results from GPT-4V.
Figure 5: We explore how countries are distributed on a color spectrum by first calculating a global average RGB vector and then defining deltas along each axes aggregated at the country level. Takeaway: We find interesting associations - Greece is strongly correlated with blue, Burkina Faso with red.
This process results in many unique cultural artifacts associated with a single country. To narrow this down to the salient artifacts that occur relatively more frequently for one country but less for others (i.e. stereotypical associations), we follow an approach similar to Jha et al. (2024): we compute the term frequencies of each artifact for each country and also compute document frequency as the number of times an artifact occurs across all countries. The final tf-idf score for a given artifact for a particular country is obtained by multiplying term frequency and the inverse of the document frequency. We then use the distribution of these scores by calculating the mean and standard deviation across all scores, and then perform a qualitative evaluation of both outliers and scores that fit the global distribution.
Figure 6: We identify more than 18,000 unique cultural artifacts across all countries as part of our second task, and then filter them to find salient ones. This figure shows strongest correlated artifacts for 5 randomly picked countries. More examples in Appendix C.4.13.
Color Associations for Countries
We calculate the mean RGB vector for each synthetic Dalle Street image, and then calculate the average to define a global mean vector. We repeat this process at the country level to obtain a mean vector for each country. Then, we find the distance of each country from the global mean and split across the three dimensions of the vector to find colors that are more likely to be associated with some countries. One of the challenges of this approach is that we only consider a limited number of color axes which may not be sufficient to capture finer signals about more complicated color associations, similar to McCarthy et al. (2019). We leave this for future work.
Counting the Number of People
In our initial observations, we noticed that DALL-E 3 tends to generate different population densities in images from different countries for otherwise identical prompts. To explore this, our object detection prompt also reports the count of people for each image, split in two buckets (between $1$ and $10$ people, or more than $10$ people). We further divide the first bucket into two smaller buckets based on the exact count (less than $5$ people, or between $5$ and $10$ people), to get $3$ buckets overall. We process this output to consider all possible variations of these terms (for example, people, person, man, woman, etc) and aggregate statistics at the country level to find the distribution of population densities in generated images.
4.2 Results
We analyse all the DALL-E 3 generated images using GPT-4V to discover hidden associations that these LMMs makes between countries and cultural artifacts.
Selected Examples of Associations within Standard Deviation
Figure 6 (and Figure C.4.13) illustrates some of the randomly sampled interesting cultural artifacts we discover, and Figure 4 shows the distribution of the tf-idf scores we calculate for each country-artifact pair. Artifacts that fall outside the range as denoted by the standard deviation lines from the mean, would be very strongly associated with a particular country and are thus very important to more closely look at, to consider if they are stereotypical (especially negatively) or not. We include statistics for the number of artifacts we identify along with more examples of interesting associations in Appendix C.4 Table LABEL:tab:artifact-interesting.
Color associations
We find that models tend to not only associate particular cultural artifacts with countries, but also colors. Figure 5 (and Figure C.4.18) shows a subset of all the countries in the Dalle Street images versus the delta values in the RGB vector components from the global average vector. Once again, for some countries, these delta values fall well outside the range denoted by the standard deviation lines from the mean and these represent strong presence of that color component in images for that country. For example, Greece is very strongly linked to the colors blue and green, whereas Indonesia and Netherlands are associated with red. An explanation for why this may be happening may be inferred from inspecting the generated images. Images of Greece typically feature a lot of sea, blue decorations, and blue patterns (usually stereotypical) whereas Netherlands is famous for its red tulip fields, which is also an artifact association that we are able to obtain from our previous analysis.
Figure 7: We explore people count associations being made by DALL-E 3 and GPT-4V, and show some selected countries where generated images in Dalle Street have more than $10$ detected people. Takeaway: African countries typically fall into high-person-count buckets in our experiments.
Figure 8: Our CultureAdapt pipeline identifies the fine-grained elements of an image that should be modified for cultural adaptation. Note that any in-painting technique can be used. Compare CultureAdapt with DALL-E 3, which generates a completely different image. More examples of edits in Appendix C.4.14.
People-Count Associations
We parse the information obtained about population density per DALL-E 3 image from our object detection prompt for GPT-4V, to get $3$ buckets overall (between $1$ and $5$ people, between $5$ and $10$ people or more than $10$ people). Figure 7 (and  C.4.19) shows a sample of countries for each bucket versus the number of images for that country which fall into that bucket.
A general trend is that images usually fall into the two extreme buckets, and we have very few occurrences of the middle bucket. Cameroon gets the highest number of occurrences in the bucket of more than $10$ people counts, despite other countries like China and India having much higher population densities in the real world. In general, African countries fall into the high-person-count buckets, while European ones fall largely in the few-people one. This might reflect the models’ perception of societies as largely individualistic or collectivist. That said, as there might be errors by GPT-4V in the person-counting process in the images –which we have not validated except from looking at a handful of examples– we recommend taking these results with a grain of salt, and leave this further exploration for future work.
5 Cultural Adaptation (Task 3)
We propose a highly composable method, CultureAdapt, for fine-grained editing that aims to use the discovered cultural artifacts from Task $2$, and adapts a given source image to a target culture. This has many potential downstream applications. For example, in generative AI applications, if the generated content is detected to not follow the specified country label, then our pipeline can be used to update it in near real-time. Another useful application of our pipeline would be as a source of data augmentation for pre-training corpora. We can use the pipeline to create culturally augmented data and also include translations to create a parallel corpus across languages and cultures which can then be used for continued pre-training to mitigate identified issues.
5.1 Methods: CultureAdapt
Recent work on cultural translation Khanuja et al. (2024); Li and Zhang (2023); Fung et al. (2024) defines different approaches for adapting images or text from one culture to another. We propose a simple automated and highly composable pipeline that can extract culture specific entities from an image, find the most salient ones, ground them with bounding boxes, and then modify only these objects without affecting the rest of the image. After this cultural adaptation process, we use CLIPScore Hessel et al. (2021) as a reference-free metric to measure the image-country compatibility using CLIP Radford et al. (2021) (treating the name of the country as the caption).
Our pipeline is built with fully customizable components. For example, we currently use GPT-4V vision-preview for open-vocab object detection, Grounding DINO Liu et al. (2023b) for grounding objects in bounding boxes and Stable Diffusion $2$ inpainting Rombach et al. (2021) for updating these grounded object masks. But, we could also use something like Tag2Text Huang et al. (2023) or RAM Zhang et al. (2023) to generate captions for an image, and then extract tags from these captions corresponding to objects, which can then be passed into Grounded SAM Ren et al. (2024) to obtain bounding boxes and segmentation masks, which will finally be used by any inpainting model like Stable Diffusion 3 AI (2024) or MimicBrush Chen et al. (2024). We demonstrate the feasibility of our approach.
Evaluation
Let image $I_{1}$ originally correspond to country $C_{1}$. We use CultureAdapt to translate it to $I_{2}$ for country $C_{2}$. The CLIPScore of an image-country pair is denoted as $S(I,C)$. We define two delta values from this as follows:
$\Delta_{1}$ $=S(I_{2},C_{1})-S(I_{1},C_{1})$
$\Delta_{2}$ $=S(I_{2},C_{2})-S(I_{1},C_{2})$
If $\Delta_{1}<0$ and $\Delta_{2}>0$, then it would imply that after cultural adaptation, the image is closer to the target country $C_{2}$ than the source country $C_{1}$. This is exactly the success criterion for our task.
5.2 Results
We show examples of using CultureAdapt in Figure 8, where we start with a Dalle Street image for the concept of front door in the country Greece. Then we culturally translate these to $3$ other countries, China, India and USA. For all of these images shown, $\Delta_{1}<0$ and $\Delta_{2}>0$ (values in Appendix C.4.3), implying that quantitatively we have measured a culture change for all 3 cases. We include more examples of CultureAdapt in Appendix C.4.14.
We further contrast these with images generated for the same concept, country pairs from DALL-E 3 to highlight differences in the two approaches. Images generated from DALL-E 3 may be more representative of the target country as per CLIPScore but they are stylistically very different from the original source image (i.e lack object consistency before and after the cultural translation process), thus making them less usable for different applications where this property is necessary.
6 Human Studies
We choose a small subset of our data, about $300$ images randomly sampled, and perform a couple of human studies with $14$ people. Here we present highlight findings due to space, with more details and annotator demographics in Appendix C.3.
Country Representation in Dalle Street
We examine whether annotators consider specific images to be stereotypical representations of a country. Figure C.4.20(a) shows that most participants agree or strongly agree, with only $2$ out of $304$ images receiving disagreement. When unsure, participants tend to neither agree nor disagree about the appropriateness of an image, . This study indicates that our generated images effectively represent country-level associations. Our participants also mark this information with visual cues (Table C.LABEL:tab:artifact-human), from explicit (e.g. flags) to implicit (e.g. color choice, nuanced cultural artifacts). Based on this analysis, we develop our framework for cultural artifact extraction and cultural adaptation using Dalle Street images.
Human baseline for Cultural Awareness
We look at accuracy metrics from $5$ different perspectives. For our study, we allowed users to select a minimum of $1$ label and a maximum of $5$ per image. These labels can be at the country level, or at the continent level or subcontinent level. So, we can evaluate exact match accuracy first at the country level, then at the subregion level and finally at the continent level. In addition, we also consider the case of union (correct answer at one of the three levels is present in the human provided labels), and the case of intersection (correct answer at all three levels are present in the human provided labels). Figure C.4.20(b) shows that on average participants have a low accuracy at the country level, but nearly double performance every time we go up a geographical level, which is along the lines of what we expect.
7 Conclusion
This study addresses the critical need for cultural awareness in Large Multimodal Models (LMMs) by introducing a comprehensive framework to evaluate and enhance their cultural competence. We create a large-scale, culturally diverse dataset of $9,935$ images across $67$ countries and $10$ concept classes, facilitating robust evaluation of LMMs on cultural awareness tasks. Further, we introduce a cultural artifact extraction task to identify over $18,000$ cultural artifacts that co-occur frequently with these countries, revealing significant insights into the implicit cultural associations encoded in these models. We also propose a cultural adaptation task and a highly composable pipeline CultureAdapt to adapt images across cultural contexts with fine-grained edits. Overall, this work emphasizes the importance of developing culturally sensitive AI systems and provides a foundational benchmark for future research towards improvement in cultural representation.
Limitations
Coverage
Our work focuses on a limited number of concept classes across a small portion of all countries in the world. While we are able to cover most geographic regions and sub-regions with this data, it is important to ensure wider coverage over time to ensure that no culture is left behind.
Open Vocabulary Object Detector
Our pipeline for automated cultural adaptation is only as good as its components. The first and most crucial component of meaningful object tag extraction is approached using GPT-4V in our work, but open-source alternatives like RAM and Tag2Text are slowly catching up to cover more wider variety of classes.
Stable Diffusion Biases
While we are using Stable Diffusion for inpainting masked objects only, our approach does not guarantee that the object generated from the diffusion process will not reflect any harmful stereotypes. Our CLIPScore based metric for measuring country deltas is also preliminary and limited by the strength of the underlying CLIP model embedddings and its inherent biases as well.
Mitigation Strategies
An important aspect of studying differences in culture awareness of LMMs is understanding how to develop approaches to mitigate them. This may involve looking more closely at pre-training corpora and developing efficient techniques for data augmentation for the same, so that secondary approaches like prompting in local languages can extract more aligned responses. While exploring all of these is beyond the current scope of our paper, we present a survey of the most relevant recent work that attempts to do so.
Attributing model behavior to specific training examples
Methods like Park et al. (2023) aim to better understand issues with training data by assigning scores to data points which cause the model to behave in a certain way. This is useful as a first step to trace training data that reflects potentially harmful stereotypes from the list of automatically identified cultural artifacts that we identify. Recent work Hall et al. (2023) has looked at the LAION corpus Schuhmann et al. (2022) to determine that geographical regions may only co-occur with more stereotypical image captions. For example, a Greek person having dinner in Greece is less likely to caption a photo of that scene as “having food in Greece” but a tourist is more likely to mention something like “when in Greece, you must moussaka” in a caption for the same image, leading to perpetuation of stereotypes when these image-caption pairs are used for training models. Similarly, color associations that we find in this paper may simply be because images were scraped from travel websites which usually portray a more vibrant view of a country than the native viewpoint.
Data Augmentation
There are many interesting approaches to generate higher quality data to mitigate cultural differences. Dollar Street for example uses photographers as a source of collecting this data manually, whereas MaRVL uses crowdsourcing. Other automated approaches include knowledge acquisition by scraping from internet sources Fung et al. (2024), simulating conversations between LLMs Li et al. (2024b), semantic data augmentation Li et al. (2024a) and value preference aligment by collecting preferences during conversations with LLMs Kirk et al. (2024). For multimodal data augmentation, methods include approaches based on finding synonyms from semantic graphs of image captions Li and Zhang (2023) and end-to-end pipelines for image transcreation Khanuja et al. (2024). Our approach CultureAdapt aims to contribute to this area of research by automatically identifying cultural artifacts and providing a composable pipeline for finer-grained edits to these artifacts.
Prompting in local languages
Hall et al. (2023) find that images generated with non-English languages tend to struggle with prompt consistency and continue to show stereotyped representations, whereas Pouget et al. (2024) find that pre-training with global, unfiltered data before fine-tuning on English content can improve cultural understanding without sacrificing performance on popular benchmarks. These contrasting findings imply the need to dedicate more efforts to find the impact of pre-training on data that has been augmented for visual modality using pipelines such as ours along with translated image captions for a multimodal parallel corpora Etxaniz et al. (2024) to eliminate geographical stereotypes. Improving the quality of the text encoders used in LMMs is another important angle to look at in this light.
Acknowledgements
This project was generously supported by the National Science Foundation under grant IIS-2327143 and by the Microsoft Accelerate Foundation Models Research (AFMR) grant program.
Appendix A Related work
Recent spike in interest for culturally aware NLP has produced many relevant works, which has inspired different aspects of our research.
Li and Zhang (2023) uses semantic graphs to develop an annotation free data augmentation approach for augmenting cultural components of captions. The main challenge of this approach is that the cultural artifacts identified for replacement are copy-pasted into the target image, which leads to inconsistencies at the object boundaries. Similarly, Khanuja et al. (2024) define the image transcreation task formally in the context of culture, and develop pipelines for cultural translation. The edited images from these pipelines tend to vary in structure, and may sometimes look very different from the source image. Our proposed CultureAdapt pipeline overcomes both of these challenges, by using automatically generated object labels grounded with bounding boxes to create precise semantic masks which are then filled in with a diffusion based model in an inpainting setting.
Qiu et al. (2024) calculates co-occurence statistiscs for features extracted from images and Jha et al. (2024) assigns importance scores to attributes to determine those which co-occur more frequently for an identity group over others. We explore both of these ideas in our work to understand which cultural artifacts are more likely to co-occur for a given country. Pouget et al. (2024) and Hall et al. (2023) both explore the two real-world datasets, MaRVL and Dollar Street, specifically developing reliable metrics to measure geo-localization and object consistency across regions. Liu et al. (2024) defines a taxonomy for culturally aware and adapted NLP which we borrow terminology from, whereas Adilazuarda et al. (2024) provides a survey of the current state of research in this sub-field and also what is yet to be done.
Other works also explore downstream tasks, for example, Ignat et al. (2024) looks at cross-cultural inspiration detection in social media data, Seaborn et al. (2024) develops voice assistants with cultural sensitivity and Srinivasan et al. (2024) improves cultural diversity of search results.
Appendix B Dalle Street generation prompt
Dalle Street Generation Prompt A typical scene of {category} in {country}, culturally accurate and detailed.
Figure 9: We use a simple prompt that includes information about the concept class and the target country using a template, to generate our large scale dataset of DALL-E 3 images.
Appendix C Appendix
C.1 Dataset details
Dollar Street concept classes (10)
car, cups, mugs and glasses, family snapshots, front door, home, kitchen, plate of food, social drink, wall decoration and wardrobe.
Dollar Street countries (63)
South Africa, Serbia, Indonesia, Brazil, Kenya, India, Nigeria, France, Kazakhstan, United States, Philippines, Mexico, Sri Lanka, Netherlands, Thailand, Colombia, Pakistan, China, Russia, Egypt, Iran, United Kingdom, Romania, Spain, Turkey, Ukraine, Italy, Czech Republic, Denmark, Ethiopia, Jordan, Burundi, Burkina Faso, Malawi, Somalia, Zimbabwe, Haiti, Cote d’Ivoire, Myanmar, Papua New Guinea, Liberia, Cambodia, Bangladesh, Rwanda, Nepal, Palestine, Tunisia, Cameroon, Bolivia, Ghana, Vietnam, Guatemala, Mongolia, South Korea, Kyrgyzstan, Lebanon, Tanzania, Switzerland, Sweden, Canada, Peru, Austria and Togo.
Concept classes (10) for Dalle Street
car, cups, mugs and glasses, family snapshots, front door, home, kitchen, plate of food, social drink, wall decoration and wardrobe.
Countries in our Dalle Street (67)
Austria, Bangladesh, Bolivia, Brazil, Bulgaria, Burkina Faso, Burundi, Cambodia, Cameroon, Canada, China, Colombia, Cote d’Ivoire, Czech Republic, Denmark, Egypt, Ethiopia, France, Ghana, Greece, Guatemala, Haiti, India, Indonesia, Iran, Italy, Jordan, Kazakhstan, Kenya, Kyrgyzstan, Latvia, Lebanon, Liberia, Lithuania, Malawi, Mexico, Mongolia, Myanmar, Nepal, Netherlands, Nigeria, Pakistan, Palestine, Papua New Guinea, Peru, Philippines, Romania, Russia, Rwanda, Serbia, Somalia, South Africa, South Korea, Spain, Sri Lanka, Sweden, Switzerland, Tanzania, Thailand, Togo, Tunisia, Turkey, Ukraine, United Kingdom, United States, Vietnam, Zimbabwe
MaRVL Country to Language Mappings
In the context of the MaRVL dataset, various languages are mapped to specific sub-regions based on the countries where these languages are predominantly spoken. The mapping is as follows:
  • "id": The language code for Indonesian, which is primarily spoken in Indonesia, corresponds to the South-eastern Asia sub-region.
  • "sw": The language code for Swahili, used in countries such as Tanzania, Kenya, and Rwanda, is mapped to the Eastern Africa sub-region.
  • "ta": The language code for Tamil, spoken in India and Sri Lanka, is associated with the Southern Asia sub-region.
  • "tr": The language code for Turkish, which is the official language of Turkey, falls under the Western Asia sub-region.
  • "zh": The language code for Chinese, predominantly spoken in China, is linked to the Eastern Asia sub-region.
Table 2: Dataset Statistics
Sub-region Eastern Africa Eastern Asia South-eastern Asia Southern Asia Western Asia Caribbean Central America Central Asia Eastern Europe Melanesia
MaRVL 875 1107 1091 924 917 - - - - -
Dollar Street 310 313 578 839 128 56 12 20 136 14
DALL-E 3 Images 1052 438 840 742 600 160 176 280 741 147
Total 2237 1858 2509 2505 1645 216 188 300 877 161
Sub-region Middle Africa Northern Africa Northern America Northern Europe South America Southern Africa Southern Europe Western Africa Western Europe Total
MaRVL - - - - - - - - - 4914
Dollar Street 107 81 317 51 447 60 223 262 183 4137
DALL-E 3 Images 303 289 465 736 605 139 740 888 594 9935
Total 410 370 782 787 1052 199 963 1150 777 18986
C.2 Prompt details
Prompt used for Object Detection with GPT-4V
We use a detailed prompt for GPT-4V to extract objects, colors and counts from images generated with DALL-E 3.
GPT-4V Object Detection Prompt Give me a json output of the items you see in this image in both the foreground and background. Output the objects as a JSON with two fields: ’relevant_objects’ for objects pertinent to the image category concept and ’other_objects’ for all additional detected objects. Be as specific as possible. Within each field, for each detected object, include sub-fields describing object attributes like color, count, and anything else that is appropriate. For example, for buildings describe the architectural style in a sentence, for people describe clothing and headgear (if multiple colors and headgears are present, include the top three), for food items describe the exact type of food and include a brief recipe description, for pictures of rooms include objects in the background like mountains outside a window or paintings on the wall portraying something specific like a landmark or a particular type of scenery. For the counts of items, if the number of items is less than 10, give me exact numbers otherwise say more than 10.
Prompt used for processing generated objects
We use a detailed prompt for GPT-4 to process the dictionaries generated with the previous prompt into a simplified list along with some parsing rules to ensure correctness of the data structure.
GPT-4 Processing Object Detection Outputs You will be provided a dictionary of items for the country {country} and concept {concept}. Summarize the dictionary into a list of comma-separated list of items with their respective colors. For example, [red apple, blue car, green tree, house with a red roof and tinted glasses]. Strictly follow the output format requested. The dictionary is as follows:
C.3 Human Study - Interface and Annotators
Annotator Demographics
All annotators are demographically located in the USA and are between 25-40 years old. Roughly 40% identify as female, and the rest as male. In terms of an education background, 40% of the annotators are graduate students, whereas the rest includes working professionals from different backgrounds and also computer science faculty. In total, we have 14 annotators, recruited from different computer science labs at an university and also from a diverse set of social connections for this study. All the annotators have agreed to consent for using this data for research purposes. Our study qualifies for exemption from IRB as no PII is involved.
General Instructions for Human Study
We use https://labelstud.io Tkachenko et al. (2020–2022) to perform our human study. For each annotator, we create 2 tabs corresponding to the 2 studies, and ask them to solve them in numerical order to avoid potential biases from seeing true labels first from the second study. Time taken to complete the first study is usually 2 hours, and the time taken to complete the second study is typically 30 minutes. All annotators will be compensated for their time with a \$20 gift card upon completion of the task.
Task 1 Instructions
For every image, you have to make atleast 1 guess for the geographical region label, along with atleast 1 corresponding clue. If you are not sure what the clue is, add a question mark symbol at the end of it - example, headgear? bread?). Do not reverse image search or look anything up. Answer only using your knowledge or instinct. You can try guessing sub-region/region if you are not sure about country. You can use the knowledge of the fact that the image is generated by an AI conditioned on the provided prompt above the image. Note that you do not have to be correct!
Figure 10: Annotation Interface for Study 1
Once the label is done, you need to add at least one bounding box somewhere in the image (it can be very specific and small or very broad or even the entire image) and then label that bounding box as either a clue or a stereotypical clue or a confusing element and then add a text description for the bounding box from the interface on the right (for example, a bounding box for a basket of baguettes can be a clue for France and the text description may be either something specific like “baguette” or something generic like “bread?”). The difference between stereotypical clue and regular clue is that stereotypical would be something like baguettes for France or “naan” for India or specific clothing styles for some country whereas a regular clue is something that you are using to make your guess but you don’t know enough about your guess to know what stereotypical clues might be associated with it, for example, sand for island countries.
Figure 11: Annotation Interface for Study 2
Task 2 Instructions
For every image, select a single rating for “appropriateness” $=$ “this image is one of the possible (stereo)typical representation of the mentioned category for the mentioned country”. Then select at least one clue corresponding to your rating, similar to the first study.
C.4 Additional Results
$\Delta_{1}$ $\Delta_{2}$
Greece $\rightarrow$ China $-8.48$ $5.07$
Greece $\rightarrow$ India $-3.47$ $0.11$
Greece $\rightarrow$ USA $-6.13$ $5.81$
Table 3: Delta values for cultural adaptation from source country to target country. A successful cultural adaptation is indicated by $\Delta_{1}<0$ and $\Delta_{2}>0$.
Performance of Human Study participants on the Cultural Awareness Task
We include anonymized performance statistics of each of our annotators to show differences in performance at the individual data point level for countries, subregions and continents.
User Country Level Subregion Level Continent Level
User 1 46.53 70.14 90.97
User 2 16.67 31.94 78.47
User 3 11.19 31.47 70.63
User 4 32.81 50.78 72.66
User 5 21.13 51.41 75.35
User 6 10.87 32.61 71.01
User 7 28.67 64.34 84.62
User 8 22.14 43.57 69.29
User 9 7.09 34.04 70.92
User 10 26.43 62.86 83.57
User 11 20.71 50.71 87.14
User 12 4.86 18.05 75.69
User 13 3.57 17.85 72.14
User 14 6.47 37.41 75.53
Table 4: User accuracies across country, subregion and continent levels, rounded to two decimal places. At the country level, accuracy varies from 46% to about 4%, so the subregion level accuracies are a more reliable indicator of performance even for humans. Continent is the most generic label, and has very high accuracies from all participants.
How many artifacts did we identify using GPT-4V
Table 5 shows the counts across each of the 67 countries for the number of artifacts identified. adj implies unique artifacts are a combination of words and adjectives that appear before it to quantify count or color. no_adj implies only the raw words identified. Figure 4 includes a distribution of the TD-IDF scores for these (country, artifact) pairs and high scores that lie outside the range as given by the mean and the standard deviation of the distribution (i.e larger than 3.01 or smaller than 0.47) would imply strongly correlated artifacts for a given country, and there exists 4019 such items from this data. Further human filtering can be done to remove common words like table or mailbox or dresses to find unique and interesting associations like pretzels in Austria, zinnias in Bolivia and many more, some of which we report in Table LABEL:tab:artifact-interesting.
Table 5: Artifact Statistics
Austria Bangladesh Bolivia Brazil Bulgaria Burkina Faso Burundi Cambodia Cameroon Canada
adj 248 283 279 264 264 288 292 276 251 292
no_adj 154 148 141 157 138 153 141 161 133 170
China Colombia Cote d’Ivoire Czech Republic Denmark Egypt Ethiopia France Ghana Greece
adj 275 268 270 276 278 276 252 278 265 270
no_adj 124 134 146 160 168 152 139 157 137 151
Guatemala Haiti India Indonesia Iran Italy Jordan Kazakhstan Kenya Kyrgyzstan
adj 272 301 264 312 246 262 280 257 270 269
no_adj 163 160 147 172 123 135 158 139 149 147
Latvia Lebanon Liberia Lithuania Malawi Mexico Mongolia Myanmar Nepal Netherlands
adj 269 244 276 267 281 266 282 292 290 260
no_adj 156 133 163 159 152 133 156 155 156 146
Nigeria Pakistan Palestine Papua New Guinea Peru Philippines Romania Russia Rwanda Serbia
adj 279 254 270 279 265 276 272 244 288 250
no_adj 152 144 137 141 137 152 149 144 161 143
Somalia South Africa South Korea Spain Sri Lanka Sweden Switzerland Tanzania Thailand Togo
adj 291 272 279 263 250 260 265 272 273 279
no_adj 165 165 144 149 150 157 155 145 154 140
Tunisia Turkey Ukraine United Kingdom United States Vietnam Zimbabwe Total
adj 249 254 267 281 273 291 311 18212
no_adj 133 132 148 181 162 163 166 10035
More examples of edits using CultureAdapt
We include more examples of edits across different concept classes and source-target pairs using our CultureAdapt pipeline in Figure 14. As can be seen, the pipeline is only constrained by the two bottlenecks of object detection and diffusion based inpainting, which sometimes may detect objects incorrectly or not generate consistent images of human faces for example.
Figure 12: We normalize income data from Dollar Street into region specific quartiles and plot corresponding accuracies for LLaVA.
Figure 13: We identify more than 18,000 unique cultural artifacts across all countries as part of our second task, and then filter them to find salient ones. This figure shows strongest correlated artifacts for 20 randomly picked countries.
Note that these associations are extracted from LMM generations and may not always be accurate.
Figure 14: We show examples of edits made using our CultureAdapt pipeline across 4 different concept classes and 12 pairs of unique source, target combinations to illustrate both cases where our pipeline excels and also where it is limited by the parts it is composed of. For all of these edits, our metric success criteria of $\Delta_{1}<0$ and $\Delta_{2}>0$ is satisfied.
Figure 15: Confusion matrices for GPT-4V and LLaVA on the cultural awareness task for Dalle Street images.
Figure 16: Confusion matrices for GPT-4V and LLaVA on the cultural awareness task for Dollar Street images.
Figure 17: Confusion matrices for GPT-4V and LLaVA on the cultural awareness task for MaRVL images.
Figure 18: We explore how countries are distributed on a color spectrum by first calculating a global average RGB vector for Dalle Street images and then defining deltas along each axes aggregated at the country level. Takeaway: We find interesting associations - Greece is strongly correlated with blue, Burkina Faso with red.
Figure 19: Here, we look at buckets of people counts in Dalle Street images aggregated at the country level, each of the subplots representing one bucket. Takeaway: Counts of people in images may not always accurately reflect population densities of the corresponding countries to scale.
(a) The counts of answers obtained from our human study about the appropriateness of generated images shows that mostly people agree or are neutral, with only 2 disagreements overall.
(b) We measure accuracy for the cultural awareness task at the country, subregion and region level and find increasing performance in this order across our annotator pool.
Figure 20: (a) Appropriateness of generated images as perceived by human study participants. (b) Accuracy of cultural awareness at various geographical levels across our annotator pool for a subset of randomly sampled Dalle Street images.
Table 6: Interesting associations and their explanations for various countries. Note that these associations are extracted from LMM generations and may not always be accurate.
Country Interesting Associations and Explanations
Austria Dirndl: A traditional dress worn in Austria and parts of Germany.
Pretzel: A type of baked bread product, often associated with German-speaking countries.
Lederhosen: Traditional leather shorts worn by men in the Alpine regions.
Bangladesh Lungi: A traditional garment worn by men, usually a wraparound skirt.
Kurti: A traditional garment worn by women, often paired with leggings or a skirt.
Harmonium: A musical instrument commonly used in South Asian music.
Bolivia Zinnias: A type of flower native to the region, known for its bright colors and significance in local celebrations.
Llama: A domesticated South American camelid, significant in Bolivian culture.
Chullos: Knitted hats, typically with ear flaps, that are traditional to the Andes.
Brazil Bikini: Associated with the famous beaches of Brazil.
Lychee: A tropical fruit found in Brazil.
Samba: A Brazilian music genre and dance style.
Bulgaria Spanakopita: A savory pastry filled with spinach and feta cheese.
Moussaka: A layered dish with eggplant, potatoes, and minced meat.
Terracotta: Refers to clay-based unglazed or glazed ceramic.
Cameroon Kaftans: A type of long robe worn in many African countries.
Fufu: A dough-like food made from cassava or yams.
Savanna: A type of ecosystem common in Cameroon, characterized by grassland with scattered trees.
Canada Poutine: A dish consisting of fries topped with cheese curds and gravy.
Moose: A large mammal found in Canada.
Snowmobile: A vehicle designed for travel on snow, common in Canadian winters.
China Changshan: A traditional Chinese garment for men.
Baozi: A type of Chinese steamed bun with fillings.
Lion Dance: A traditional dance in Chinese culture performed during the Lunar New Year and other cultural events.
Ethiopia Injera: A sourdough flatbread and a staple food in Ethiopia.
Wat: A traditional Ethiopian stew.
Shawl: Often worn by Ethiopian women as part of traditional attire.
France Camembert: A famous French cheese.
Baguette: A long, thin loaf of French bread.
Beret: A soft, round, flat-crowned hat associated with French culture.
Germany Oktoberfest: An annual beer festival and cultural event in Munich.
Bratwurst: A type of German sausage.
Dirndl: Traditional dress worn by women during Oktoberfest and other occasions.
Greece Toga: A garment worn in ancient Greece.
Dolma: A dish made of grape leaves stuffed with rice or meat.
Moussaka: A layered dish with eggplant, meat, and béchamel sauce.
India Sari: A traditional garment worn by women.
Lassi: A yogurt-based drink.
Rangoli: A form of art created on the floor using colored rice, sand, or flower petals.
Japan Kimono: A traditional Japanese garment.
Sushi: A popular Japanese dish.
Tatami: A type of mat used as a flooring material in traditional Japanese rooms.
Mexico Sombrero: A wide-brimmed hat traditionally worn in Mexico.
Tacos: A traditional Mexican dish.
Guacamole: A Mexican avocado-based dip or spread.
Morocco Tagine: A North African dish named after the earthenware pot in which it is cooked.
Kaftan: A long robe worn in Morocco.
Mint Tea: A popular beverage in Morocco, often served as a welcoming gesture.
Nepal Topi: A traditional hat worn in Nepal.
Himalayas: The mountain range running across Nepal.
Dal Bhat: A traditional Nepalese dish consisting of lentils and rice.
Peru Chullo: A traditional hat with earflaps.
Llama: A significant animal in Peruvian culture.
Ponchos: Traditional clothing made from wool.
Thailand Tuk-tuk: A common form of transportation in Thailand.
Pad Thai: A popular Thai noodle dish.
Elephant: An animal deeply ingrained in Thai culture and symbolism.
Togo Kente Cloth: A traditional fabric made of silk and cotton, known for its vibrant colors and patterns.
Yam Festival: A major cultural festival celebrating the harvest of yams.
Agbadza Dance: A traditional dance performed during festivals and ceremonies.
Tunisia Shisha: A popular water pipe used for smoking flavored tobacco.
Harissa: A spicy chili paste that is a staple in Tunisian cuisine.
Mosaic Art: Intricate and colorful tile art that is significant in Tunisian culture.
Turkey Evil Eye: A common talisman believed to protect against negative energy.
Baklava: A sweet pastry made of layers of filo filled with nuts and honey.
Whirling Dervishes: A religious dance performed by Sufi practitioners.
Ukraine Pysanky: Traditional Ukrainian Easter eggs decorated with intricate designs.
Borscht: A beet soup that is a key part of Ukrainian cuisine.
Vyshyvanka: Traditional Ukrainian embroidered shirts.
United Kingdom Afternoon Tea: A British tradition involving tea and a variety of snacks.
Red Telephone Box: Iconic public telephone booths found throughout the UK.
Fish and Chips: A classic British dish of battered fish and fried potatoes.
United States Route 66: A historic highway symbolizing the American road trip.
Thanksgiving: A national holiday celebrating the harvest and other blessings.
Statue of Liberty: A symbol of freedom and democracy in the US.
Vietnam Ao Dai: A traditional Vietnamese dress for women.
Pho: A Vietnamese noodle soup that is a staple dish.
Conical Hat (Non La): A traditional hat made of bamboo and palm leaves.
Zimbabwe Mbira: A traditional musical instrument also known as the thumb piano.
Great Zimbabwe: The ruins of an ancient city, significant in Zimbabwean history.
Victoria Falls: One of the largest and most famous waterfalls in the world, located on the border between Zimbabwe and Zambia.
Table 7: Cultural artifacts for various countries based on human annotations. Note that these artifacts are based on subjective perceptions of our human annotators and may not be completely accurate always.
Country Cultural Artifacts
Austria beer, sausage, dirndl
Bangladesh rice, saree, fish
Bolivia colorful clothes, poncho, hats
Brazil brazilian flag, tropical fruit, colorful pottery
Bulgaria clothing, rugs, door
Burkina Faso dry, black people, straw basket
Burundi rice, beans, bananas
Cambodia buddhism, buddhist art, clothing
Cameroon african people, bananas, beans
Canada maple leaf, canadian flag, poutine
China characters, chinese food, lanterns
Colombia coffee, rice, avocado
Cote d’Ivoire black people, dry, african outfit
Czech Republic beer, dress, czech
Denmark danish flag, beer, windmill
Egypt hieroglyphs, egyptian art, islamic clothing
Ethiopia coffee, colors, clay pots
France baguette, cheese, wine
Ghana black people, african necklaces, clothing
Greece blue and white, sea, olives
Guatemala mayan art, tortilla, beans
Haiti black people, rice, beans
India naan, curry, sari
Indonesia buddhism, rice, clothing
Iran islamic art, kebab, persian rug
Italy pizza, pasta, wine
Jordan clothing, arabic, islamic art
Kazakhstan clothing, houses, islamic art
Kenya african people, african art, corn
Kyrgyzstan clothing, islamic art, rugs
Latvia clothing, beer, bread
Lebanon arabic clothing, hummus, bread
Liberia rice, black people, palm trees
Lithuania clothing, food, beer
Malawi hut, corn, black people
Mexico sombrero, tequila, tortilla
Mongolia yurt, dumplings, clothing
Myanmar buddhist art, rice, pagoda
Nepal buddhist elements, hindu elements, rice
Netherlands windmill, cheese, dutch clothing
Nigeria rice, yams, african clothing
Pakistan clothing, curry, sombrero
Palestine arabic art, hummus, bread
Papua New Guinea black people, tropical fruit, coconut
Peru inca clothing, machu picchu, andes mountains
Philippines rice, tropical vegetation, cooking
Romania clothing, sheep, ceramic pots
Russia fur hat, warm clothes, vodka
Rwanda african art, beans, dark-skinned people
Serbia clothing, beer, sausages
Somalia islamic art, banana, rice
South Africa african art, corn, hat
South Korea korean characters, kimchi, korean dress
Spain flamenco, paella, bull fighting
Sri Lanka buddhist art, buddhist symbols, spicy food
Sweden northern european clothing, fish, snowy landscape
Switzerland alps, swiss cheese, chocolate
Tanzania african art, rice, meat
Thailand buddhist art, thai food, clothing
Togo african clothing, cloth patterns, wood carvings
Tunisia arabic art, couscous, arched doorways
Turkey turkish coffee, rugs, kebabs
Ukraine clothing patterns, flower designs, ukrainian food
United Kingdom pubs, fish and chips, tea
United States american flag, burgers, jeans
Vietnam conical hats, pho, pagodas
Zimbabwe thatched huts, african clothing, animal carvings
C.5 LLM hyperparameters
We discuss the generation settings we used for our experiments, and also the associated costs and hardware.
Generation settings
  • DALL-E 3 images were generated for vivid and natural settings for standard quality and size $1024$ x $1024$
  • GPT-4 and GPT-4V generations were obtained for temperature $=0.7$, top_p $=0.95$, no frequency or presence penalty, no stopping condition other than the maximum number of tokens to generate, max_tokens $=300$.
  • LLaVA generations were obtained for temperature $=1.0$ and top_p $=1.0$, no penalties, and max_tokens $=128$. The reason for using a slightly higher temperature and top_p is to have more consistent outputs. In our initial experiments, LLaVA did not perform as well in terms of following instructions at the same temperatue setting as GPT-4V.
  • For Grounding DINO, we use ShilongLiu/GroundingDINO from Hugging Face and set both box and text thresholds to 0.25 for the grounded box generations.
  • For Stable Diffusion, we use stabilityai/stable-diffusion-2-inpainting from Hugging Face, and replace the autoencoder with stabilityai/sd-vae-ft-mse. We also use a DPMSolverMultistepScheduler for speeding up the generation process. We add "intricate details. 4k. high resolution. high quality." to the end of our prompt to get high quality images.
Computation budget
  • We spent about \$$\$800$ in total for DALL-E 3 generations. This was funded by a grant from Microsoft Azure.
  • We spent about \$$\$700$ in total for GPT-4V vision-preview and GPT-4 turbo inference and across all experiments.
  • For experiments with LLaVA, Stable Diffusion and Grounding DINO, we used a single instance of a Multi-Instance A100 GPU with 40GB of GPU memory, 3/7 fraction of Streaming Multiprocessors, 2 NVIDIA Decoder hardware units, 4/8 L2 cache size, and 1 node.
  • Total emissions for API based models are estimated to be 25 kgCO2 eq, of which 100 percents were directly offset by the cloud provider. Total emissions for our on-premise GPU usage is estimated to be less than 5 kgCO2 eq. Estimations were conducted using the MachineLearning Impact calculator Lacoste et al. (2019).

### v2

@@
 Abstract
-In this work, we present a comprehensive three-phase study to examine (1) the effectiveness of large multimodal models (LMMs) in recognizing cultural contexts; (2) the accuracy of their representations of diverse cultures; and (3) their ability to adapt content across cultural boundaries.
-We first introduce Dalle Street, a large-scale dataset generated by DALL-E 3 and validated by humans, containing $9,935$ images of $67$ countries and $10$ concept classes.
-We reveal disparities in cultural understanding at the sub-region level with both open-weight (LLaVA) and closed-source (GPT-4V) models on Dalle Street and other existing benchmarks.
-Next, we assess models’ deeper culture understanding by an artifact extraction task and identify over $18,000$ artifacts associated with different countries.
-Finally, we propose a highly composable pipeline, CultureAdapt, to adapt images from culture to culture.
+We present a comprehensive three-phase study to examine ($1$) the cultural understanding of Large Multimodal Models (LMMs) by introducing Dalle Street, a large-scale dataset generated by DALL-E 3 and validated by humans, containing $9{,}935$ images of $67$ countries and $10$ concept classes; ($2$) the underlying implicit and potentially stereotypical cultural associations with a cultural artifact extraction task; and ($3$) an approach to adapt cultural representation in an image based on extracted associations using a modular pipeline, CultureAdapt.
+We find disparities in cultural understanding at geographic sub-region levels with both open-source (LLaVA) and closed-source (GPT-4V) models on Dalle Street and other existing benchmarks, which we try to understand using over $18{,}000$ artifacts that we identify in association to different countries.
 Our findings reveal a nuanced picture of the cultural competence of LMMs, highlighting the need to develop culture-aware systems.
@@
 1 Introduction
-Figure 1: We introduce a large scale balanced dataset for measuring cultural awareness, a task for extracting implicit cultural artifacts from images and a highly composable pipeline for adapting images to different cultural contexts with fine-grained edits.
+Figure 1: We introduce a large-scale dataset for measuring cultural awareness, an artifact extraction task for implicit cultural associations, and a modular pipeline for culturally adapting images with fine-grained edits.
 Culture is hard to define and has always been so.
@@
 (2024).
-As large multimodal models (LMMs; Liu et al., 2023a) intersect more with human life, the need for them to comprehend and respect cultural nuances is crucial.
-Research in this area explores model alignment with human values, measures of cultural awareness, and methods for cultural adaptation, where content that may be considered as accurately representing a culture or a country, maybe stereotypically, is modified to represent a different one.
-The challenge of assessing the capability of understanding and leveraging cultural knowledge in LMMs, given their multimodal nature – processing both text and image data – is significant.
+As large multimodal models (LMMs) intersect more with human life, the need for them to comprehend and respect cultural nuances is crucial.
+Research in this area focuses on model alignment with human values, assessing cultural awareness, and exploring cultural adaptation; the goal is to modify content that represents one culture or country, often stereotypically, to reflect a different one, to suit audiences from different cultural backgrounds better.
+The challenge of assessing the capability of understanding and leveraging cultural knowledge in LMMs is significant.
 Prior research has primarily investigated LMMs for cultural awareness [^2] We maintain that LMMs do not inherently possess human values but that their outputs may display cultural knowledge.
-by examining their performance on culture-related tasks such as region classification from images Basu et al.
+by examining their performance on tasks such as region classification from images Basu et al.
 (2023); Yin et al.
@@
 (2024).
-However, these tasks may not fully determine whether LMMs are responding to cultural cues encoded within their training data or merely identifying superficial culture-associated features.
-Furthermore, there is a lack of in-depth analysis of how LMMs utilize cultural knowledge during these tasks and whether they exhibit consistent understanding across various cultural contexts.
-To address these gaps, We propose a comprehensive three-phase study to deepen our understanding of cultural awareness in LMMs.
-First, we develop a large-scale dataset for assessing cultural awareness as measured by the ability of LMMs to recognize and differentiate between cultures, with countries as proxies, in a task setting similar to GeoGuessr Geoguessr (2024).
-Next, we delve deeper into cultural understanding by introducing an artifact extraction task designed to identify implicit cultural artifacts that LMMs use to distinguish between cultures.
-Finally, based on the insights gained from the artifact extraction task, we propose a cultural adaptation task combining multiple generative models in an end-to-end pipeline to adapt images from one cultural context to another.
-This pipeline enhances our understanding of LMMs’ adaptability and promises an effective data augmentation technique to improve cultural awareness in LMMs.
+However, these tasks do not determine whether LMMs are responding to cultural cues encoded within their training data or merely identifying superficial cultural associations.
+To address these gaps, first, we develop a new large-scale dataset to assess cultural awareness as measured by the ability of LMMs to recognize and differentiate between cultures, with countries as proxies, in a task setting similar to GeoGuessr Geoguessr (2024).
+Next, we introduce a task designed to identify implicit associations between cultures and artifacts Liu et al.
+(2024) that LMMs use to distinguish between cultures.
+Finally, we propose a cultural adaptation framework combining multiple generative models in an end-to-end pipeline to adapt images from one cultural context to another by modifying the underlying implicit associations.
 Our main contributions are as follows:
-  • We introduce, Dalle Street, a collection of $9,935$ images generated by DALL-E 3, covering $67$ countries and $10$ cultural concept classes.
-Compared to datasets like Dollar Street Rojas et al.
-(2022), Dalle Street includes more images from underrepresented geographic regions.
-We validate the appropriateness of country representations with a human study and will release this dataset publicly.
-  • We measure how well both open-weight and closed-source LMMs perform on Dalle Street based on the cultural knowledge encoded in them to understand disparities in performance at the geographic sub-region level for a diverse group of concepts and countries.
-We include a human study baseline to understand where current models may perform better than humans.
-We also compare LMM performance on two image datasets collected from different sources.
-  • We introduce a task for cultural artifact extraction to identify implicit associations.
-We further develop a method to filter these effectively to discover interesting associations that frequently co-occur for each country.
-  • We propose a highly composable end-to-end pipeline, CultureAdapt, using a chain of LMMs to identify a country for a given image, extract culturally relevant artifacts, ground them with bounding boxes, and use diffusion-based inpainting to adapt to a target culture.
-We show examples of this pipeline and include a CLIPScore-based metric to validate its performance.
+  • Dataset: We introduce Dalle Street, a collection of $9{,}935$ images generated by DALL-E 3, covering $67$ countries and $10$ cultural concept classes, with more images from underrepresented geographic regions compared to datasets like Dollar Street Rojas et al.
+(2022).
+  • Benchmark: We measure how well humans and multimodal large language models (open- and closed-source) can identify countries for images in Dalle Street and two other datasets (Dollar Street, MaRVL) to study disparities in performance at the geographic subregion level for a diverse group of concepts and countries.
+  • Task: We introduce a task for identifying implicit associations by extracting cultural artifacts from images and filtering them to discover associations that frequently co-occur for each country.
+  • Framework: We propose a modular end-to-end pipeline, CultureAdapt (Figure 10), to adapt an image to a target culture by updating identified implicit cultural associations in it using diffusion-based inpainting.
+We evaluate results by introducing a CLIPScore-based metric.
 2 Data
-We use three different datasets covering a wide variety of cultural concepts, economic ranges, and different sources: (a) Dalle Street: synthetic, DALL-E 3 (OpenAI, 2024) generated; (b) Dollar Street: natural, collected photographs; and (c) MaRVL: scraped from the internet under native speaker guidance.
-We include count statistics for each of the datasets in Table C.2.
-Overall, our study uses roughly $20$k images across $3$ datasets, covering $19$ geographical sub-regions and $67$ countries from the United Nations geoscheme UN (2024).
-Each image is accompanied by a true label for country.
+We study around $20$k images from three datasets (data statistics in Table 2) covering a wide variety of cultural concepts, economic ranges, and data sources: (a) Dalle Street: synthetic, DALL-E 3 OpenAI (2024) generated; (b) Dollar Street: natural, collected photographs; and (c) MaRVL: web-scraped under native speaker guidance.
+All the datasets are available under CC BY-SA $4.0$ license.
 Dalle Street
-We consider the same $10$ categories and $19$ geographical regions as in Dollar Street and include $67$ countries.
-[^3] The four extra countries corresponding to new additions to our annotator pool over the timeline of this project.
-Given these parameters, we generate high resolution images of size $1024$x1024102410241024 from DALL-E 3 OpenAI (2024) for 2 settings - vivid (for generating hyper-real and dramatic images) and natural (for more realistic images).
-To do so, we follow a templatic prompting approach (prompt shown in Appendix Figure 9) and then sample at least $10$ images for each combination of country and concept class.
-However, due to global content policy filters imposed on the API calls and limited budget, we end up with $9,935$ images (as opposed to the expected $13,400$).
-We will release this generated dataset under CC BY-SA 4.0.
-We perform a qualitative study, where we ask annotators to determine the “appropriateness” of the generated image on a Likert scale from $1$ (Strongly Disagree) to $5$ (Strongly Agree), where we define appropriateness as the case where the image shown to the annotator is one of the (possibly stereotypical) representations for the given combination of country and category.
+We use $10$ concept classes (car, family snapshots, front door, home, kitchen, plate of food, cups/mugs/glasses, social drink, wall decoration, and wardrobe), $19$ geographical regions, and $67$ countries (Section A.1), similar to Dollar Street.
+We generate $1024\times 1024$ images with DALL-E 3 OpenAI (2024) in two styles - vivid (hyper-realistic) and natural (realistic), prompting with a template (Figure 14) that specifies the concept class and target country.
+At least $10$ images are sampled per country-concept combination, yielding $9{,}935$ images after filtering out content policy violations from API calls.
+A qualitative study on a randomly sampled subset of around $300$ generated images and $14$ participants (Table 1(a)) shows most annotators agree that the images reflect stereotypical country representations, with less than $1\%$ of images receiving strong disagreement.
+When unsure, participants tend to neither agree nor disagree about the “appropriateness” of an image.
+Our annotators also mark visual cues in these images, including both explicit (e.g., flags) and implicit (e.g., color schemes) cultural artifacts.
+This feedback motivates our artifact extraction task and its use in our cultural adaptation framework.
 Dollar Street Rojas et al. (2022)
-This is a dataset of photos of objects and scenes collected by professional and volunteer photographers which specifically focuses on poor and remote environments, containing approximately $200$ concept classes and covering $63$ countries.
-We filter it for images that do not contain multiple labels, classes which do not cover images for all regions, and classes with subjective naming.
+This is a dataset of photos of objects and scenes collected by professional and volunteer photographers.
+We filter it for images that do not contain multiple labels, classes that do not cover images for all regions, and classes with subjective naming.
 Then we refer to our group of annotators from diverse backgrounds to choose the top $10$ categories by the method of collaborative labeling Chang et al.
-(2017), where effectively we simplify our selection of object classes by choosing the ones which all annotators universally agree on as being a relevant dimension for testing cultural awareness.
-The final data subset that we end up with thus includes $63$ countries, $19$ geographical regions, $4,137$ images, $10$ concept classes ( car, cups, mugs and glasses, family snapshots, front door, home, kitchen, plate of food, social drink, wall decoration, and wardrobe).
-This dataset is available under CC BY-SA 4.0.
+(2017), where we simplify our selection of object classes by choosing the ones which all annotators universally agree on as being a relevant dimension for testing cultural awareness.
+Our data from this source includes $4{,}137$ images from $63$ countries, $19$ geographical regions, across $10$ concept classes.
 MaRVL Liu et al. (2021)
-The primary task for this dataset involves validation of statements about image pairs for five languages: Indonesian, Mandarin Chinese, Swahili, Tamil, and Turkish.
-We use the images and assign country and region labels to create our task setting for cultural awareness.
-[^4] MaRVL language to region mapping in Appendix C.1 This dataset gives us $4,914$ images curated by native speakers of each of these languages, across $5$ distinct geographical regions and it is distributed under CC BY 4.0 license for research purposes.
+This is originally a dataset for validation of statements about image pairs curated by native speakers in five languages: Indonesian, Mandarin Chinese, Swahili, Tamil, and Turkish.
+We assign country and region labels (based on corresponding languages) to get $4{,}914$ images across $5$ geographical regions.
 3 Cultural Awareness (Task 1)
-To measure cultural awareness of LMMs, we compare performance on two existing culturally diverse benchmarks and also on Dalle Street.
-Later on, we also compare this with the baseline from our human study to understand where LMMs stand at this task in comparison to humans.
-Overall, we find performance varies across sub-regions but both LLaVA and GPT-4V perform nearly equally well.
+We compare performances of humans, LLaVA and GPT-4V on Dalle Street, Dollar Street and MaRVL in terms of their ability to predict the country given an image.
+Overall, we find performances vary across sub-regions, but both LLaVA and GPT-4V perform better than humans.
+Appropriateness Category Percentage (%)
+Agree $40.79$
+Neither Agree nor Disagree $34.54$
+Strongly Agree $19.41$
+Disagree $4.61$
+Strongly Disagree $0.66$
+(a) Results from our human study on appropriateness for generated images show that most participants agree or are neutral, with less than $1$% expressing strong disagreement.
+Geographical Level Accuracy (%)
+Country Level $22.16$
+Subregion Level $47.63$
+Continent Level $77.77$
+Union Accuracy $78.03$
+Intersection Accuracy $21.91$
+(b) Accuracy for the cultural awareness task improves from country to subregion to region level.
+Table 1: (a) Perceived appropriateness of generated images by human participants. (b) Cultural awareness accuracy at different geographical levels for a subset of Dalle Street images.
 3.1 Methods
-Given an input image, we prompt two LMMs, LLaVA-NeXT (Liu et al., 2023a), an open-weight model, and GPT-4V vision-preview (OpenAI et al., 2023), a closed-source model, in a zero-shot setting by asking an open-ended question without any possible answer choices: predict the geographical region represented in the image, as per the United Nations geoscheme.
-We use this geoscheme for three reasons.
-First, the refusal rate of models is a lot higher when asked about specific country labels.
-Second, the geoscheme information is already present on English Wikipedia, which is usually included in pre-training corpora for most LLMs.
-And third, sticking to a geoscheme allows us to parse open-ended generations in a structured format.
-A potential challenge with this approach is that models may not incorporate the most recent changes in the geoscheme, leading to finer grained errors which would be hard to trace (for example, the Wikipedia page for the geoscheme has been updated more than $20$ times since the cut-off date for the training data of vision-preview, including some changes to island groups like Channel Islands).
-We focus on countries which have traditionally been consistently classified without changes to geographic region to avoid these inconsistencies.
+Given an input image, we prompt LLaVA-NeXT Liu et al.
+(2023a) and GPT-4V vision-preview OpenAI et al.
+(2023) in a zero-shot generative setting by asking an open-ended question without providing answer choices: Predict the geographical region represented in the image, as per the United Nations geoscheme UN (2024).
+We use this geoscheme for three reasons: ($1$) models have a higher refusal rate when queried with specific country labels; ($2$) the geoscheme is included in most LLM pre-training data (English Wikipedia); and ($3$) it enables structured parsing of open-ended generations.
+We focus on countries with stable geographic classifications to prevent errors due to geoscheme updates.
+Evaluation Metrics
+We process generated text to map it to one of the geographical sub-regions or a policy violation case and then compare it with true labels by mapping country information to geographical regions, which gives us classification accuracy as a quantitative metric for measuring success.
+Since this is a typical classification problem, we also inspect the confusion matrix to locate sub-regions with more errors.
+Economic disparities
+For Dollar Street, we also have data available for the monthly income of the family corresponding to each image.
+We use this information to understand differences in performance across economic groups by looking at region-specific normalized income quartiles.
 GPT-4V LLaVA
@@
 MaRVL $\mathbf{41.59}$ $19.14$
-Table 1: Overall, LLaVA performs as good as or better than GPT-4V on two out of three datasets. Human accuracy on a subset of Dalle Street is $42.63$.
-Metric: Accuracy
-We process generated text to map it to one of the geographical sub-regions or to a policy violation case, and then compare with true labels by mapping country information to geographical regions, which gives us classification accuracy as a quantitative metric for measuring success.
-Specifically, we inspect the confusion matrix to better understand on which regions the models make mistakes.
-Economic disparities
+Figure 2: LLaVA matches or outperforms GPT-4V on two of three datasets. Human accuracy on a Dalle Street subset is 47.63%.
+Figure 3: Confusion matrices for GPT-4V on the cultural awareness task for Dalle Street images.
+Accurate responses match the true subregion.
+Special labels include Invalid (no match or incomplete) and ResponsibleAI (policy violation).
+Takeaway: The model performs well, with a strong leading diagonal and $100\%$ accuracy for Western Asia (which covers Iran, Jordan, Lebanon, Oman, Palestine, Turkey).
+Human Baseline
+$14$ annotators label images at the country, subregion, or continent level, with $1$ to $5$ guesses per image, accounting for varying familiarity with different regions.
+We first evaluate exact match accuracy at the country, then subregion, and finally continent levels.
+We also consider two cases: union (the correct answer appears at any level) and intersection (the correct answer appears at all levels).
+Table 1(b) shows that while country-level accuracy is low, it nearly doubles at each broader geographic level (Table 3).
+3.2 Results
+We find similar trends across datasets for both models, with some variations across subregions.
+Figure 4: We normalize Dollar Street income data into region-specific quartiles and plot accuracies for GPT-4V.
+Takeaway: Lower income quartiles (Q1111, Q2222) show higher accuracy in Africa and Asia, while higher quartiles (Q3333, Q4444) perform better in the Americas.
+In Europe, accuracy is similar across all quartiles.
+Overall comparison
+LLaVA performs as good as GPT-4V on Dollar Street and outperforms it significantly on the Dalle Street images (Table 3).
+This indicates that LLaVA may have implicitly learned stereotypical associations between regions and concepts because the Dalle Street images include such associations (Section 4).
+However, on the MaRVL data, LLaVA performs about as good as random guessing.
+This may be because MaRVL covers specific indigenous concepts that the model may not have seen before.
+Subregion Level Analysis
+GPT-4V performs well on Dalle Street, with a strong leading diagonal indicating many correct predictions (Figure 3).
+However, it often provides no answer due to content policy violations.
+Notably, both models accurately predict all Western Asian images (Figure 23).
+LLaVA tends to default to South America for incorrect answers, while GPT-4V defaults to policy violations.
+Similar trends are observed in other datasets (Figures 24, 25).
+Economic Disparity
+Using income data from Dollar Street, we group results by normalized income quartiles across Africa, Asia, Americas, Europe, and Oceania (GPT-4V - Figure 4, LLaVA - Figure 22).
+Performance is better for lower-income quartiles in Africa and Asia, while it improves with higher-income groups for the Americas.
+This might indicate that the model defaults to associating Africa with poorer contexts and America with wealthier ones.
+For Europe, performance remains consistent across all quartiles.
+4 Extracting Implicit Associations of Cultures and Artifacts (Task 2)
+We propose to extract cultural artifacts (material items) from the generated images to identify the implicit associations the models may use for Task $1$.
+We find associations that are usually stereotypical (and not truly representative) for the relevant countries.
+This provides a better understanding of the models, enabling us to develop our approach for Task $3$ for cultural adaptation of images.
+Figure 5: We score each artifact based on its likelihood of co-occurrence for a country.
+Scores outside the mean and standard deviation range (red) indicate frequent co-occurrences, representing implicit (potentially stereotypical) associations.
+4.1 Methods
+We use GPT-4V vision-preview for open vocabulary object detection Zareian et al.
+(2021), using a detailed prompt (Figure 16) to extract information about concept classes in Dalle Street images, including descriptions, color, [^3] This refers to object and person appearance, not race.
+and person count.
+GPT-4V’s strong instruction-following capabilities result in nearly perfect JSON outputs, which we lightly post-process and summarize using GPT-$4$ turbo (prompt in Figure 17).
+This process yields many unique associations for each country.
+In our initial experiments, GPT-4V significantly outperformed LLaVA, hence we report GPT-4V results.
+Salient associations
+To identify salient artifacts that appear more frequently in one country than others (potentially stereotypical associations), we follow an approach similar to  Jha et al.
+(2024): compute the term frequencies of each artifact for each country and also compute document frequency as the number of times an artifact occurs across all countries, to calculate a tf-idf score by multiplying term frequency and the inverse of the document frequency.
+We then perform a qualitative evaluation of outliers from the distribution of these scores.
+Figure 6: We identify more than $18{,}000$ unique cultural artifacts across all countries as part of our second task and then filter them to find salient ones (Table 4).
+This figure shows the strongest correlated artifacts for $5$ randomly picked countries.
+We include more such examples of country-object associations in the Appendix (Figure 26).
+Evaluations
 […]
-Human Study - GeoGuessr
-We develop an annotation interface similar to Geoguessr (2024) and ask our pool of annotators to label images at either the country or the region level or at continent level.
-We ask for a maximum of $5$ labels and a minimum of $1$ per image, to adjust for different familiarity levels with multiple randomly sampled regions/cultures.
-Appendix C.3 includes detailed descriptions of annotator demographics, interface and instructions provided to annotators.
+Figure 7: Human validation of a random sample of generated artifacts shows (1) low hallucination (high presence scores) and (2) more than a random fraction of artifacts usually display cultural relevance.
+Color Associations for Countries
+We calculate the mean RGB vector for each Dalle Street image, then average them to get a global mean vector.
+We repeat this process at the country level and measure each country’s distance from the global mean, identifying colors more strongly associated with specific countries across the three RGB dimensions.
+Counting the Number of People
+We observed that DALL-E 3 generates varying population densities across countries for identical prompts.
+To explore this, we use an object detection prompt (Figure 16) to count people in each image, split into three buckets: less than $5$, $5$ to $10$, and more than $10$ people.
+We process related terms (e.g., people, person, man, woman) and aggregate country-level statistics to analyze population density distributions in the generated images.
+Annotators validate a random sample of these counts, finding general consistency, though they often do not reflect real-world population densities accurately.
+Figure 8: We explore how countries are distributed on a color spectrum by first calculating a global average RGB vector and then defining deltas along each axes aggregated at the country level.
+Takeaway: We find interesting associations - Greece is strongly correlated with blue, and the Netherlands with red/orange.
 4.2 Results
-Contrary to our expectations, LLaVA is not far off from GPT-4V in terms of cultural awareness, as measured in our experimental settings for this task.
-Overall comparison
-We show the aggregated results for our Cultural Awareness Task in Table 1.
-LLaVA performs as good as GPT-4V on Dollar Street and even outperforms it significantly on the Dalle Street images.
-This is a potential indication that LLaVA might be more prone to forming stereotypical associations between regions and concepts, because the Dalle Street images include such associations to distinguish different geographical regions.
-As we discuss later, it might be possible to trace this down to the training data.
-However, we note that on the MaRVL data, LLaVA performs about as good as random guessing, specifically getting incorrect predictions for nearly all of the images corresponding to Swahili and Turkish.
-Our hypothesis to explain this is based on the fact that the other two datasets have similar distributions about concepts and countries, whereas MaRVL is much more diverse and covers fewer countries but many indigenous concepts, not all of which the model may have seen during training.
-Figure 2: Confusion matrices for GPT-4V on the cultural awareness task for Dalle Street images.
-Accurate responses match the true subregion exactly.
-Special labels include Invalid (no match to region labels or incomplete) and ResponsibleAI (no region label due to policy violation).
-Takeaway: The model shows balanced performance, indicated by a strong leading diagonal.
-Notably, the model achieves $100$% accuracy for Western Asia images, covering Iran, Jordan, Lebanon, Oman, Palestine, and Turkey.
-Subregion level analysis
-Figure  2 (and Figure  C.4.15) indicates that both GPT-4V and LLaVA have strong leading diagonals, which is an indicator that they get many predictions correct.
-GPT-4V has many cases where it does not provide any answer due to violations of the content policy filter.
-Instead of trying to find workarounds for the filter, we show it as is, because this would represent the experience of any typical user of the system.
-Very interestingly, both GPT-4V and LLaVA gets all of the answers from Western Asia correct.
-In general, LLaVA makes a lot of mistakes for different regions, but defaults to South America as the incorrect answer for these cases, whereas for GPT-4V the same translates to the ResponsibleAI policy filter output as the incorrect answer.
-We include results for the other two datasets in the Appendix (Figure C.4.16 and  C.4.17 ), and find similar trends.
-Figure 3: We normalize income data from Dollar Street into region specific quartiles and plot corresponding accuracies for GPT-4V.
-Takeaway: Lower income quartiles (blue, orange bars) are more accurate in Africa and Asia, but higher income quartiles (red, green) are more accurate in Americas; in Europe all quartiles have similar accuracies.
-Economic disparity - GPT-4V
-We utilize the monthly income information available as part of the Dollar Street data, and divide the images into groups based on normalized income quartiles for each broad geographical region - Africa, Asia, Americas, Europe and Oceania (see Figure 3).
-Lower income quartiles show better performance for Africa and Asia, whereas for America performance improves for higher income groups.
-This discrepancy reveals the latent biases of the model, which seem to associate Africa with poorer contexts and America with richer ones.
-For Europe, performance is nearly similar for all quartiles.
-4 Automated Extraction of Implicit Cultural Artifacts (Task 2)
-To identify the implicit associations that the models may be using for performing Task $1$, we propose a second task to extract specific cultural artifacts from the images, and discover interesting associations, which are usually stereotypical for the relevant countries.
-Figure 4: We assign a score for each identified artifact based on how likely it is to co-occur for a given country.
-Across all such scores, we find the mean and standard deviation to define a range.
-Scores that lie outside this range would correspond to items that co-occur frequently for a given country, and would most likely represent stereotypical implicit associations.
+We analyze all the DALL-E 3 generated images to discover implicit associations between countries and cultural artifacts.
+Our method effectively surfaces implicit associations but also highlights challenges with stereotype reinforcement and demographic inaccuracies.
+Artifact Associations
+Our analysis shows that some cultural artifacts are strongly tied to specific countries, with certain artifacts exceeding one standard deviation from the mean tf-idf score (Figure 5).
+While these associations can offer cultural insights, they often reflect negative stereotypes, such as the over-representation of palm trees for tropical regions, which overlooks broader diversity.
+This suggests that while our method captures implicit associations, it also underscores the need to refine models to avoid reinforcing such stereotypes.
+Color Associations
+Models not only associate cultural artifacts with countries but also colors.
+In Figure 8, the RGB delta values for several countries in Dalle Street fall outside the standard deviation.
+For instance, Greece is strongly associated with blue and the Netherlands with red, likely due to recurring elements like blue seas in Greek images and red tulip fields in Dutch ones.
+People-Count Associations
+Most images fall into the extreme buckets (less than $5$ or more than $10$ people), with few in the middle (Figures 28,9), often misrepresenting actual population densities.
+In general, African countries tend to fall into high person-count buckets, while European ones are on the low person-count end, possibly reflecting the model’s perception of collectivist versus individualistic societies.
+5 Cultural Adaptation (Task 3)
+We propose a method to edit a given image for a target culture by modifying the detected salient implicit associations between countries and artifacts.
 5.1 Methods
-We use the open vocabulary object detection Zareian et al.
-(2021) capabilities of GPT-4V vision-preview, formulating a detailed prompt (see Appendix C.2) to extract information about the target concept class in a Dalle Street image, including text descriptions, color, [^5] Note that this corresponds to appearance of different objects and persons in images, and does not imply race.
-and person count.
-GPT-4V’s strong instruction following capabilities allow the generation of a nearly perfectly validated JSON string, which we lightly post-process, and then provide that as input to GPT-4 turbo to get a summary of the objects for any given image.
-In our initial experiments, GPT-4V outperformed LLaVA significantly in terms of following complex instructions for object detection, so we only report results from GPT-4V.
-Figure 5: We explore how countries are distributed on a color spectrum by first calculating a global average RGB vector and then defining deltas along each axes aggregated at the country level.
-Takeaway: We find interesting associations - Greece is strongly correlated with blue, Burkina Faso with red.
-This process results in many unique cultural artifacts associated with a single country.
-To narrow this down to the salient artifacts that occur relatively more frequently for one country but less for others (i.e.
-stereotypical associations), we follow an approach similar to Jha et al.
-(2024): we compute the term frequencies of each artifact for each country and also compute document frequency as the number of times an artifact occurs across all countries.
-The final tf-idf score for a given artifact for a particular country is obtained by multiplying term frequency and the inverse of the document frequency.
-We then use the distribution of these scores by calculating the mean and standard deviation across all scores, and then perform a qualitative evaluation of both outliers and scores that fit the global distribution.
-Figure 6: We identify more than 18,000 unique cultural artifacts across all countries as part of our second task, and then filter them to find salient ones.
-This figure shows strongest correlated artifacts for 5 randomly picked countries.
-More examples in Appendix C.4.13.
-Color Associations for Countries
-We calculate the mean RGB vector for each synthetic Dalle Street image, and then calculate the average to define a global mean vector.
-We repeat this process at the country level to obtain a mean vector for each country.
-Then, we find the distance of each country from the global mean and split across the three dimensions of the vector to find colors that are more likely to be associated with some countries.
-One of the challenges of this approach is that we only consider a limited number of color axes which may not be sufficient to capture finer signals about more complicated color associations, similar to McCarthy et al.
+Recent works on cultural translation Khanuja et al. (2024); Li and Zhang (2023); Fung et al. (2024) define different approaches for adapting images or text from one culture to another.
+CultureAdapt
+Our pipeline (Figure 10) uses GPT-4V for open-vocab object detection to extract implicit cultural associations from the source image.
+Next, we use Grounding DINO Liu et al.
+(2023b) to ground these objects with bounding boxes, which we convert into masks.
+We then create an inpainting prompt by adding the target country to the list of detected objects and use Stable Diffusion $2$ inpainting Rombach et al.
+(2021) to edit [^4] We acknowledge that the resulting image may represent common stereotypes because the underlying artifacts may be implicitly associated with a stereotypical view of the country.
+the image by filling in the masked pixels.
+We evaluate our method using CLIPScore Hessel et al.
+(2021) to measure image-country similarity (treating the name of the country as the caption) and cosine similarity of DINO-ViT Caron et al.
+(2021) embeddings for measuring structural preservation.
+Our choice of metrics is inspired from  Khanuja et al.
 (2024).
-We leave this for future work.
-Counting the Number of People
-In our initial observations, we noticed that DALL-E 3 tends to generate different population densities in images from different countries for otherwise identical prompts.
-To explore this, our object detection prompt also reports the count of people for each image, split in two buckets (between $1$ and $10$ people, or more than $10$ people).
-We further divide the first bucket into two smaller buckets based on the exact count (less than $5$ people, or between $5$ and $10$ people), to get $3$ buckets overall.
-We process this output to consider all possible variations of these terms (for example, people, person, man, woman, etc) and aggregate statistics at the country level to find the distribution of population densities in generated images.
-4.2 Results
-We analyse all the DALL-E 3 generated images using GPT-4V to discover hidden associations that these LMMs makes between countries and cultural artifacts.
-Selected Examples of Associations within Standard Deviation
-Figure 6 (and Figure C.4.13) illustrates some of the randomly sampled interesting cultural artifacts we discover, and Figure 4 shows the distribution of the tf-idf scores we calculate for each country-artifact pair.
-Artifacts that fall outside the range as denoted by the standard deviation lines from the mean, would be very strongly associated with a particular country and are thus very important to more closely look at, to consider if they are stereotypical (especially negatively) or not.
-We include statistics for the number of artifacts we identify along with more examples of interesting associations in Appendix C.4 Table LABEL:tab:artifact-interesting.
-Color associations
-We find that models tend to not only associate particular cultural artifacts with countries, but also colors.
-Figure 5 (and Figure C.4.18) shows a subset of all the countries in the Dalle Street images versus the delta values in the RGB vector components from the global average vector.
-Once again, for some countries, these delta values fall well outside the range denoted by the standard deviation lines from the mean and these represent strong presence of that color component in images for that country.
-For example, Greece is very strongly linked to the colors blue and green, whereas Indonesia and Netherlands are associated with red.
-An explanation for why this may be happening may be inferred from inspecting the generated images.
-Images of Greece typically feature a lot of sea, blue decorations, and blue patterns (usually stereotypical) whereas Netherlands is famous for its red tulip fields, which is also an artifact association that we are able to obtain from our previous analysis.
+Modularity
+Our pipeline is modular, allowing components to be easily swapped to improve performance over time.
+For example, we could use Tag2Text Huang et al.
+(2023) or RAM Zhang et al.
+(2023) for image captioning, extract object tags, and then use Grounded SAM Ren et al.
+(2024) for bounding boxes and segmentation masks.
+These can then be passed to inpainting models like Stable Diffusion 3 AI (2024) or MimicBrush Chen et al.
+(2024).
 Figure 9: We explore people count associations being made by DALL-E 3 and GPT-4V, and show some selected countries where generated images in Dalle Street have more than $10$ detected people.
 Takeaway: African countries typically fall into high-person-count buckets in our experiments.
-Figure 8: Our CultureAdapt pipeline identifies the fine-grained elements of an image that should be modified for cultural adaptation.
-Note that any in-painting technique can be used.
-Compare CultureAdapt with DALL-E 3, which generates a completely different image.
-More examples of edits in Appendix C.4.14.
-People-Count Associations
-We parse the information obtained about population density per DALL-E 3 image from our object detection prompt for GPT-4V, to get $3$ buckets overall (between $1$ and $5$ people, between $5$ and $10$ people or more than $10$ people).
-Figure 7 (and  C.4.19) shows a sample of countries for each bucket versus the number of images for that country which fall into that bucket.
-A general trend is that images usually fall into the two extreme buckets, and we have very few occurrences of the middle bucket.
-Cameroon gets the highest number of occurrences in the bucket of more than $10$ people counts, despite other countries like China and India having much higher population densities in the real world.
-In general, African countries fall into the high-person-count buckets, while European ones fall largely in the few-people one.
-This might reflect the models’ perception of societies as largely individualistic or collectivist.
-That said, as there might be errors by GPT-4V in the person-counting process in the images –which we have not validated except from looking at a handful of examples– we recommend taking these results with a grain of salt, and leave this further exploration for future work.
-5 Cultural Adaptation (Task 3)
-We propose a highly composable method, CultureAdapt, for fine-grained editing that aims to use the discovered cultural artifacts from Task $2$, and adapts a given source image to a target culture.
-This has many potential downstream applications.
-For example, in generative AI applications, if the generated content is detected to not follow the specified country label, then our pipeline can be used to update it in near real-time.
-Another useful application of our pipeline would be as a source of data augmentation for pre-training corpora.
-We can use the pipeline to create culturally augmented data and also include translations to create a parallel corpus across languages and cultures which can then be used for continued pre-training to mitigate identified issues.
-5.1 Methods: CultureAdapt
-Recent work on cultural translation Khanuja et al.
-(2024); Li and Zhang (2023); Fung et al.
-(2024) defines different approaches for adapting images or text from one culture to another.
-We propose a simple automated and highly composable pipeline that can extract culture specific entities from an image, find the most salient ones, ground them with bounding boxes, and then modify only these objects without affecting the rest of the image.
-After this cultural adaptation process, we use CLIPScore Hessel et al.
-(2021) as a reference-free metric to measure the image-country compatibility using CLIP Radford et al.
-(2021) (treating the name of the country as the caption).
-Our pipeline is built with fully customizable components.
-For example, we currently use GPT-4V vision-preview for open-vocab object detection, Grounding DINO Liu et al.
-(2023b) for grounding objects in bounding boxes and Stable Diffusion $2$ inpainting Rombach et al.
-(2021) for updating these grounded object masks.
-But, we could also use something like Tag2Text Huang et al.
-(2023) or RAM Zhang et al.
-(2023) to generate captions for an image, and then extract tags from these captions corresponding to objects, which can then be passed into Grounded SAM Ren et al.
-(2024) to obtain bounding boxes and segmentation masks, which will finally be used by any inpainting model like Stable Diffusion 3 AI (2024) or MimicBrush Chen et al.
+Figure 10: The CultureAdapt pipeline first identifies objects in an image using GPT-4V and grounds them with bounding boxes from Grounding DINO.
+These masks, along with a prompt containing object names and the target country, are used by Stable Diffusion for inpainting and cultural adaptation.
+Takeaway: Unlike DALL-E 3 (which generates a completely new image) or other editing methods Khanuja et al.
+(2024) that struggle with cultural adaptation, CultureAdapt makes precise, meaningful edits.
+Additional examples in the Appendix (Figure 29).
+Baseline
+The closest related work is by Khanuja et al.
+(2024), which proposes three methods for image editing.
+We compare our approach with their two most relevant methods, providing qualitative examples (Figure 10) and quantitative evaluations (Table 7).
+Additionally, we conduct a qualitative study assessing human preferences for layout preservation and cultural relevance changes (Figure 12) for country pairs common to both studies.
+We also perform extensive statistical testing to compare structural similarity and editing success, using CLIPScore-based metrics, for both approaches.
+Evaluation
+Let image $I_{1}$ correspond to country $C_{1}$ and $I_{2}$ be its adaptation for country $C_{2}$. The CLIPScore for an image-country pair is denoted as $S(I,C)$. We define two deltas:
+$\Delta_{1}$ $=S(I_{2},C_{1})-S(I_{1},C_{1})$ (1)
+$\Delta_{2}$ $=S(I_{2},C_{2})-S(I_{1},C_{2})$ (2)
+If $\Delta_{1}<0$ and $\Delta_{2}>0$, it indicates successful adaptation, where $I_{2}$ is closer to $C_{2}$ than $C_{1}$.
+Our primary metric, $M_{1}$, tracks how often this condition is met.
+A secondary metric, $M_{2}$, compares $\Delta_{2}-\Delta_{1}$ to evaluate success, though it’s less ideal for cases where $\Delta_{1}$ is positive.
+We also use another metric, SSIM, to measure structural similarity between $I_{1}$ and $I_{2}$ via DINO-ViT embeddings.
+A good edit performs well across editing and similarity metrics.
+5.2 Results
+Empirical findings indicate that our method works well both qualitatively and quantitatively.
+Qualitative comparisons
+In Figure 10, we show CultureAdapt applied to a randomly selected image, for adapting from Greece to China, India, and the USA.
+By visually contrasting with results from DALL-E 3, we see that our approach preserves structural similarity better, and by comparing with the baselines e2e-instruct and cap-edit, we demonstrate our method’s ability to make meaningful edits.
+Figure 29 includes more examples.
+Comparison with baseline
+We compare CultureAdapt with cap-edit (Figure 11) for $20$ country pairs (Table 7). Both methods produce images similar to the source, but CultureAdapt edits are overall more culturally relevant.
+To test this empirically, we compare similarity scores using the Wilcoxon signed-rank test (Shapiro-Wilk: $p=5.99\times 10^{-39}$, indicating non-normality).
+cap-edit has a slightly higher mean similarity $0.97$ than CultureAdapt $0.94$, with a statistically significant difference ($p=6.02\times 10^{-215}$) for ($\alpha=0.05$).
+The bootstrapped $95$% confidence interval for the mean difference is [$0.0298$, $0.0333$], which does not include 00, supporting this result.
+For editing metrics, CultureAdapt outperformed cap-edit in a statistically significant way.
+For $M_{1}$, $54\%$ of samples met the condition versus $50\%$ for cap-edit ($p=7.43\times 10^{-5}$, McNemar’s test).
+For $M_{2}$, CultureAdapt had a higher mean score ($3.11$ vs.
+$2.68$), with the difference significant per the Wilcoxon test ($p=4.39\times 10^{-11}$), as non-normality was confirmed by the Shapiro-Wilk test ($p=1.23\times 10^{-16}$).
+The bootstrapped $95\%$ confidence interval for $M_{2}$ differences supports our conclusion ([$-0.560$, $-0.307$]).
+In a human study of $100$ images, $3$ participants rated both methods equally preferable on structure preservation and cultural relevance (Figure 12).
+Figure 11: Our method performs better than other approaches in terms of editing metrics while still maintaining comparable structural similarities.
+Figure 12: When asked to compare editing outputs from CultureAdapt to the existing baseline cap-edit, our annotators have nearly similar preferences for both methods in terms of (1) the ability of the methods to maintain layout similarity and (2) cultural relevance of the edited image to the target country.
+Error Analysis
+We identify two common error modes (Figure 13): ($1$) multiple similar objects to be edited in the source image lead to masks covering a significant portion of the image, often in an overlapping manner, resulting in major changes as the diffusion process needs to inpaint from scratch, and ($2$) when editing realistic DALL-E 3 images, the edited objects sometimes lose realism.
+Replacing Grounding DINO with segmentation models (e.g., Grounded SAM) helps mitigate the first issue, but the second remains an open problem, as noted by others Hall et al.
 (2023).
-We demonstrate the feasibility of our approach.
-Evaluation
-Let image $I_{1}$ originally correspond to country $C_{1}$.
-We use CultureAdapt to translate it to $I_{2}$ for country $C_{2}$.
-The CLIPScore of an image-country pair is denoted as $S(I,C)$.
-We define two delta values from this as follows:
-$\Delta_{1}$ $=S(I_{2},C_{1})-S(I_{1},C_{1})$
-$\Delta_{2}$ $=S(I_{2},C_{2})-S(I_{1},C_{2})$
-If $\Delta_{1}<0$ and $\Delta_{2}>0$, then it would imply that after cultural adaptation, the image is closer to the target country $C_{2}$ than the source country $C_{1}$.
-This is exactly the success criterion for our task.
-5.2 Results
-We show examples of using CultureAdapt in Figure 8, where we start with a Dalle Street image for the concept of front door in the country Greece.
-Then we culturally translate these to $3$ other countries, China, India and USA.
-For all of these images shown, $\Delta_{1}<0$ and $\Delta_{2}>0$ (values in Appendix C.4.3), implying that quantitatively we have measured a culture change for all 3 cases.
-We include more examples of CultureAdapt in Appendix C.4.14.
-We further contrast these with images generated for the same concept, country pairs from DALL-E 3 to highlight differences in the two approaches.
-Images generated from DALL-E 3 may be more representative of the target country as per CLIPScore but they are stylistically very different from the original source image (i.e lack object consistency before and after the cultural translation process), thus making them less usable for different applications where this property is necessary.
-6 Human Studies
-We choose a small subset of our data, about $300$ images randomly sampled, and perform a couple of human studies with $14$ people.
-Here we present highlight findings due to space, with more details and annotator demographics in Appendix C.3.
-Country Representation in Dalle Street
-We examine whether annotators consider specific images to be stereotypical representations of a country.
-Figure C.4.20(a) shows that most participants agree or strongly agree, with only $2$ out of $304$ images receiving disagreement.
-When unsure, participants tend to neither agree nor disagree about the appropriateness of an image, .
-This study indicates that our generated images effectively represent country-level associations.
-Our participants also mark this information with visual cues (Table C.LABEL:tab:artifact-human), from explicit (e.g.
-flags) to implicit (e.g.
-color choice, nuanced cultural artifacts).
-Based on this analysis, we develop our framework for cultural artifact extraction and cultural adaptation using Dalle Street images.
-Human baseline for Cultural Awareness
-We look at accuracy metrics from $5$ different perspectives.
-For our study, we allowed users to select a minimum of $1$ label and a maximum of $5$ per image.
-These labels can be at the country level, or at the continent level or subcontinent level.
-So, we can evaluate exact match accuracy first at the country level, then at the subregion level and finally at the continent level.
-In addition, we also consider the case of union (correct answer at one of the three levels is present in the human provided labels), and the case of intersection (correct answer at all three levels are present in the human provided labels).
-Figure C.4.20(b) shows that on average participants have a low accuracy at the country level, but nearly double performance every time we go up a geographical level, which is along the lines of what we expect.
-7 Conclusion
-This study addresses the critical need for cultural awareness in Large Multimodal Models (LMMs) by introducing a comprehensive framework to evaluate and enhance their cultural competence.
-We create a large-scale, culturally diverse dataset of $9,935$ images across $67$ countries and $10$ concept classes, facilitating robust evaluation of LMMs on cultural awareness tasks.
-Further, we introduce a cultural artifact extraction task to identify over $18,000$ cultural artifacts that co-occur frequently with these countries, revealing significant insights into the implicit cultural associations encoded in these models.
-We also propose a cultural adaptation task and a highly composable pipeline CultureAdapt to adapt images across cultural contexts with fine-grained edits.
-Overall, this work emphasizes the importance of developing culturally sensitive AI systems and provides a foundational benchmark for future research towards improvement in cultural representation.
-Limitations
-Coverage
-Our work focuses on a limited number of concept classes across a small portion of all countries in the world.
-While we are able to cover most geographic regions and sub-regions with this data, it is important to ensure wider coverage over time to ensure that no culture is left behind.
-Open Vocabulary Object Detector
-Our pipeline for automated cultural adaptation is only as good as its components.
-The first and most crucial component of meaningful object tag extraction is approached using GPT-4V in our work, but open-source alternatives like RAM and Tag2Text are slowly catching up to cover more wider variety of classes.
-Stable Diffusion Biases
-While we are using Stable Diffusion for inpainting masked objects only, our approach does not guarantee that the object generated from the diffusion process will not reflect any harmful stereotypes.
-Our CLIPScore based metric for measuring country deltas is also preliminary and limited by the strength of the underlying CLIP model embedddings and its inherent biases as well.
-Mitigation Strategies
-An important aspect of studying differences in culture awareness of LMMs is understanding how to develop approaches to mitigate them.
-This may involve looking more closely at pre-training corpora and developing efficient techniques for data augmentation for the same, so that secondary approaches like prompting in local languages can extract more aligned responses.
-While exploring all of these is beyond the current scope of our paper, we present a survey of the most relevant recent work that attempts to do so.
-Attributing model behavior to specific training examples
-Methods like Park et al.
-(2023) aim to better understand issues with training data by assigning scores to data points which cause the model to behave in a certain way.
-This is useful as a first step to trace training data that reflects potentially harmful stereotypes from the list of automatically identified cultural artifacts that we identify.
-Recent work Hall et al.
-(2023) has looked at the LAION corpus Schuhmann et al.
-(2022) to determine that geographical regions may only co-occur with more stereotypical image captions.
-For example, a Greek person having dinner in Greece is less likely to caption a photo of that scene as “having food in Greece” but a tourist is more likely to mention something like “when in Greece, you must moussaka” in a caption for the same image, leading to perpetuation of stereotypes when these image-caption pairs are used for training models.
-Similarly, color associations that we find in this paper may simply be because images were scraped from travel websites which usually portray a more vibrant view of a country than the native viewpoint.
-Data Augmentation
-There are many interesting approaches to generate higher quality data to mitigate cultural differences.
-Dollar Street for example uses photographers as a source of collecting this data manually, whereas MaRVL uses crowdsourcing.
-Other automated approaches include knowledge acquisition by scraping from internet sources Fung et al.
-(2024), simulating conversations between LLMs Li et al.
-(2024b), semantic data augmentation Li et al.
-(2024a) and value preference aligment by collecting preferences during conversations with LLMs Kirk et al.
-(2024).
-For multimodal data augmentation, methods include approaches based on finding synonyms from semantic graphs of image captions Li and Zhang (2023) and end-to-end pipelines for image transcreation Khanuja et al.
-(2024).
-Our approach CultureAdapt aims to contribute to this area of research by automatically identifying cultural artifacts and providing a composable pipeline for finer-grained edits to these artifacts.
-Prompting in local languages
-Hall et al.
-(2023) find that images generated with non-English languages tend to struggle with prompt consistency and continue to show stereotyped representations, whereas Pouget et al.
-(2024) find that pre-training with global, unfiltered data before fine-tuning on English content can improve cultural understanding without sacrificing performance on popular benchmarks.
-These contrasting findings imply the need to dedicate more efforts to find the impact of pre-training on data that has been augmented for visual modality using pipelines such as ours along with translated image captions for a multimodal parallel corpora Etxaniz et al.
-(2024) to eliminate geographical stereotypes.
-Improving the quality of the text encoders used in LMMs is another important angle to look at in this light.
-Acknowledgements
-This project was generously supported by the National Science Foundation under grant IIS-2327143 and by the Microsoft Accelerate Foundation Models Research (AFMR) grant program.
-Appendix A Related work
-Recent spike in interest for culturally aware NLP has produced many relevant works, which has inspired different aspects of our research.
-Li and Zhang (2023) uses semantic graphs to develop an annotation free data augmentation approach for augmenting cultural components of captions.
-The main challenge of this approach is that the cultural artifacts identified for replacement are copy-pasted into the target image, which leads to inconsistencies at the object boundaries.
+Other less obvious issues include not maintaining the correct count or orientation of objects and not generating human faces correctly, as the underlying diffusion model is trained with a privacy filter.
+Figure 13: Common error cases with our method occur when there are multiple objects in the image or when the image is generated in a realistic style with DALL-E 3.
+6 Related Work
+The growing interest in culturally aware NLP has inspired various aspects of our research.
+Li and Zhang (2023) propose a data augmentation approach using semantic graphs to enhance cultural components in captions.
+However, their method often results in inconsistencies at object boundaries when cultural artifacts are copied and pasted into images.
 Similarly, Khanuja et al.
-(2024) define the image transcreation task formally in the context of culture, and develop pipelines for cultural translation.
-The edited images from these pipelines tend to vary in structure, and may sometimes look very different from the source image.
-Our proposed CultureAdapt pipeline overcomes both of these challenges, by using automatically generated object labels grounded with bounding boxes to create precise semantic masks which are then filled in with a diffusion based model in an inpainting setting.
+(2024) formalize the task of image transcreation, but their pipelines can produce images that differ significantly from the source or not at all.
+Our CultureAdapt pipeline maintains image coherence by generating semantic masks with bounding boxes and using diffusion-based inpainting.
 Qiu et al.
-(2024) calculates co-occurence statistiscs for features extracted from images and Jha et al.
-(2024) assigns importance scores to attributes to determine those which co-occur more frequently for an identity group over others.
-We explore both of these ideas in our work to understand which cultural artifacts are more likely to co-occur for a given country.
+(2024) calculate co-occurrence statistics from image features, and Jha et al.
+(2024) assign importance scores to attributes that frequently co-occur for identity groups.
+We build on these ideas to identify salient cultural artifacts likely to co-occur for a given country.
+Liu et al.
+(2024) provide a taxonomy for culturally aware NLP, from which we adopt terminology (cultural artifacts and adaptation).
 Pouget et al.
@@
 (2023) both explore the two real-world datasets, MaRVL and Dollar Street, specifically developing reliable metrics to measure geo-localization and object consistency across regions.
-Liu et al.
-(2024) defines a taxonomy for culturally aware and adapted NLP which we borrow terminology from, whereas Adilazuarda et al.
-(2024) provides a survey of the current state of research in this sub-field and also what is yet to be done.
-Other works also explore downstream tasks, for example, Ignat et al.
-(2024) looks at cross-cultural inspiration detection in social media data, Seaborn et al.
-(2024) develops voice assistants with cultural sensitivity and Srinivasan et al.
-(2024) improves cultural diversity of search results.
-Appendix B Dalle Street generation prompt
-Dalle Street Generation Prompt A typical scene of {category} in {country}, culturally accurate and detailed.
-Figure 9: We use a simple prompt that includes information about the concept class and the target country using a template, to generate our large scale dataset of DALL-E 3 images.
-Appendix C Appendix
-C.1 Dataset details
+7 Conclusion
+This study addresses the critical need for cultural awareness in Large Multimodal Models by introducing a comprehensive framework to evaluate and enhance their cultural competence.
+We create a large-scale, culturally diverse dataset of $9{,}935$ images across $67$ countries and $10$ concept classes, facilitating benchmarking of LMMs on cultural awareness tasks.
+Further, we introduce an artifact extraction task to identify over $18{,}000$ artifacts that co-occur frequently with these countries, revealing significant insights into the implicit cultural associations encoded in these models.
+We also propose CultureAdapt, a pipeline to adapt images across cultural contexts with fine-grained edits.
+Overall, this work emphasizes the importance of developing culturally sensitive AI systems and provides a foundational benchmark for future research toward improvement in cultural representation.
+Limitations
+Our experiments use a specific version of a closed-source API model, but later versions usually offer similar or improved performance at lower costs.
+The effectiveness of our cultural adaptation approach depends on its components, particularly object tag extraction, which currently works well when using GPT-4V.
+Open-source alternatives like RAM and Tag2Text would become more viable over time for this task.
+Our method could be extended for data augmentation to improve cultural awareness in models, though this falls outside the scope of our current work.
+Additionally, exploring multilingual prompts to analyze associations across languages would be interesting but is not our focus here.
+Since we use Stable Diffusion for inpainting, we cannot fully prevent the stereotypical generations.
+These challenges point to the need for more culturally nuanced models, as discussed in recent works Park et al.
+(2023); Hall et al.
+(2023); Li et al.
+(2024).
+Acknowledgements
+This project was generously supported by the National Science Foundation under grant IIS-2327143 and by the Microsoft Accelerate Foundation Models Research (AFMR) grant program.
+Appendix A Appendix
+  1. 1.
+    Dataset details A.1
+    1. (a)
+      Dollar Street concept classes
+    2. (b)
+      Dollar Street countries
+    3. (c)
+[…]
+    4. (d)
+[…]
+    5. (e)
+      MaRVL Country to Language Mappings
+  2. 2.
+    Prompt details A.2
+    1. (a)
+      Dalle Street generation
+    2. (b)
+      Cultural Awareness classifier
+    3. (c)
+      Object Detection
+  3. 3.
+    LLM hyperparameters A.3
+    1. (a)
+      Generation settings
+    2. (b)
+      Computation budget
+  4. 4.
+    Human Study - the Annotators A.4
+    1. (a)
+      Annotator Demographics
+    2. (b)
+      How many studies do we have
+    3. (c)
+      Total number of annotations
+  5. 5.
+    Human Study - the Interfaces A.5
+    1. (a)
+      General Instructions for Study 1 and 2
+    2. (b)
+      Study 1: Verify the quality of generated images
+      1. i.
+        Task 1 Instructions
+    3. (c)
+      Study 2: Human performance on our cultural awareness benchmark
+      1. i.
+        Task 2 Instructions
+      2. ii.
+        Anonymized performance of individual Human Study participants
+    4. (d)
+      Study 3 and 4: Artifact hallucination and cultural relevance
+    5. (e)
+      Study 5: People count associations
+    6. (f)
+      Study 6 and 7: Layout preservation and cultural relevance after editing
+  6. 6.
+    Additional Results A.6
+    1. (a)
+      Task 1 - Cultural awareness
+      1. i.
+        Performance of LLaVA on income quartiles for Dollar Street
+      2. ii.
+        Confusion matrices for LLaVA and GPT-4V on Dalle Street images
+      3. iii.
+        Confusion matrices for LLaVA and GPT-4V on Dollar Street images
+      4. iv.
+        Confusion matrices for LLaVA and GPT-4V on MaRVL images
+    2. (b)
+      Task 2 - Artifacts
+      1. i.
+        Number of artifacts identified using GPT-4V
+A.1 Dataset details
 Dollar Street concept classes (10)
-car, cups, mugs and glasses, family snapshots, front door, home, kitchen, plate of food, social drink, wall decoration and wardrobe.
+car, family snapshots, front door, home, kitchen, plate of food, cups/mugs/glasses, social drink, wall decoration, and wardrobe.
 Dollar Street countries (63)
@@
 Concept classes (10) for Dalle Street
-car, cups, mugs and glasses, family snapshots, front door, home, kitchen, plate of food, social drink, wall decoration and wardrobe.
+car, family snapshots, front door, home, kitchen, plate of food, cups/mugs/glasses, social drink, wall decoration, and wardrobe.
 Countries in our Dalle Street (67)
@@
 In the context of the MaRVL dataset, various languages are mapped to specific sub-regions based on the countries where these languages are predominantly spoken. The mapping is as follows:
-  • "id": The language code for Indonesian, which is primarily spoken in Indonesia, corresponds to the South-eastern Asia sub-region.
-  • "sw": The language code for Swahili, used in countries such as Tanzania, Kenya, and Rwanda, is mapped to the Eastern Africa sub-region.
-  • "ta": The language code for Tamil, spoken in India and Sri Lanka, is associated with the Southern Asia sub-region.
-  • "tr": The language code for Turkish, which is the official language of Turkey, falls under the Western Asia sub-region.
-  • "zh": The language code for Chinese, predominantly spoken in China, is linked to the Eastern Asia sub-region.
+  • ‘‘id’’: The language code for Indonesian, which is primarily spoken in Indonesia, corresponds to the South-eastern Asia sub-region.
+  • ‘‘sw’’: The language code for Swahili, used in countries such as Tanzania, Kenya, and Rwanda, is mapped to the Eastern Africa sub-region.
+  • ‘‘ta’’: The language code for Tamil, spoken in India and Sri Lanka, is associated with the Southern Asia sub-region.
+  • ‘‘tr’’: The language code for Turkish, which is the official language of Turkey, falls under the Western Asia sub-region.
+  • ‘‘zh’’: The language code for Chinese, predominantly spoken in China, is linked to the Eastern Asia sub-region.
 Table 2: Dataset Statistics
@@
 Total 410 370 782 787 1052 199 963 1150 777 18986
-C.2 Prompt details
+A.2 Prompt details
+A.2.1 Dalle Street generation
+Prompt for data generation
+We use a simple template to prompt DALL-E 3 to generate images for a particular combination of country and category (Figure 14).
+Dalle Street Generation Prompt A typical scene of {category} in {country}, culturally accurate and detailed.
+Figure 14: We use a simple prompt that includes information about the concept class and the target country using a template, to generate our large scale dataset of DALL-E 3 images.
+A.2.2 Cultural Awareness classifier
+Prompt for classification of images
+We use a simple prompt to generate names of subregions from models when provided an image.
+Classification Prompt Strictly follow the United Nations geoscheme for subregions.
+Which geographical subregion of the United Nations geoscheme is this image from?
+Make an educated guess.
+Answer in one to three words.
+Figure 15: We use a simple prompt to classify images in the data, by generating subregion labels.
+A.2.3 Object Detection
 Prompt used for Object Detection with GPT-4V
@@
 For the counts of items, if the number of items is less than 10, give me exact numbers otherwise say more than 10.
+Figure 16: We use a detailed prompt for GPT-4V to extract objects, colors and counts from images generated with DALL-E 3.
 Prompt used for processing generated objects
@@
 The dictionary is as follows:
-C.3 Human Study - Interface and Annotators
+Figure 17: We use a detailed prompt for GPT-4 to process the dictionaries generated with the previous prompt into a simplified list along with some parsing rules to ensure correctness of the data structure.
+A.3 LLM hyperparameters
+We discuss the generation settings we used for our experiments, and also the associated costs and hardware.
+A.3.1 Generation settings
+  • DALL-E 3 images are generated for vivid and natural settings for standard quality and size $1024\times 1024$
+  • GPT-4 and GPT-4V generations are obtained for temperature $=0.7$, top_p $=0.95$, no frequency or presence penalty, no stopping condition other than the maximum number of tokens to generate, max_tokens $=300$.
+  • LLaVA generations are obtained for temperature $=1.0$ and top_p $=1.0$, no penalties, and max_tokens $=128$.
+The reason for using a slightly higher temperature and top_p is to have more consistent outputs.
+In our initial experiments, LLaVA did not perform as well in terms of following instructions at the same temperatue setting as GPT-4V.
+  • For Grounding DINO, we use ShilongLiu/GroundingDINO from Hugging Face and set both box and text thresholds to 0.25 for the grounded box generations.
+  • For Stable Diffusion, we use stabilityai/stable-diffusion-2-inpainting from Hugging Face, and replace the autoencoder with stabilityai/sd-vae-ft-mse.
+We also use a DPMSolverMultistepScheduler for speeding up the generation process.
+We add ‘‘intricate details.
+4k.
+high resolution.
+high quality.’’ to the end of our prompt to get high quality images.
+A.3.2 Computation budget
+  • We spent about \$$\$800$ in total for DALL-E 3 generations. This was funded by a grant from Microsoft Azure.
+  • We spent about \$$\$700$ in total for GPT-4V vision-preview and GPT-4 turbo inference and across all experiments.
+  • For experiments with LLaVA, Stable Diffusion and Grounding DINO, we used a single instance of a Multi-Instance A100 GPU with 40GB of GPU memory, 3/7 fraction of Streaming Multiprocessors, 2 NVIDIA Decoder hardware units, 4/8 L2 cache size, and 1 node.
+  • Total emissions for API based models are estimated to be 25 kgCO2 eq, of which 100 percents were directly offset by the cloud provider.
+Total emissions for our on-premise GPU usage is estimated to be less than 10 kgCO2 eq.
+Estimations are conducted using the MachineLearning Impact calculator Lacoste et al.
+(2019).
+A.4 Human Study - the Annotators
 Annotator Demographics
-All annotators are demographically located in the USA and are between 25-40 years old.
+All annotators have different demographic backgrounds (but are physically located in the USA currently) and are between 25-40 years old.
+Together, they are native to or have resided in more than countries and all 4 major regions from our dataset, and thus represent a strong sample of opinions.
 Roughly 40% identify as female, and the rest as male.
@@
 Our study qualifies for exemption from IRB as no PII is involved.
-General Instructions for Human Study
+How many studies do we have
+  1. 1.
+    Human study to verify the quality of generated images
+  2. 2.
+    Human study to understand performance on our benchmark
+  3. 3.
+    Human validation for checking if artifacts are hallucinated
+  4. 4.
+    Human study to check if artifacts are culturally relevant
+  5. 5.
+    Human study for verifying people count associations
+  6. 6.
+    Human study for layout preservation after editing using CultureAdapt
+  7. 7.
+    Human study for cultural relevance after editing using CultureAdapt
+Total number of annotations
+To validate our dataset and human baseline on cultural awareness, we have $14$ people, each annotating $300$ samples, resulting in $4200$ annotations for one study and $8400$ in total.
+For each of the other three studies (excluding the studies for edits), we have three people annotating $100$ samples each for a study, resulting in $900$ annotations in total.
+Finally, for the two editing-related studies, we have three people each looking at three images per sample (source image, edited image using our method, edited image using another method) for $100$ samples, resulting in $900$ annotations per study or $1800$ in total.
+Combining all this, we have $8200+900+1800=10900$ annotation items to back the comprehensive quantitative evaluations we perform for each study in our paper.
+A.5 Human Study - the Interfaces
+General Instructions for Study 1, 2
 We use https://labelstud.io Tkachenko et al.
 (2020–2022) to perform our human study.
-For each annotator, we create 2 tabs corresponding to the 2 studies, and ask them to solve them in numerical order to avoid potential biases from seeing true labels first from the second study.
+For each annotator, we create 2 tabs corresponding to the 2 studies, and ask them to solve them in numerical order to avoid getting influenced from seeing true labels first from the second study.
 Time taken to complete the first study is usually 2 hours, and the time taken to complete the second study is typically 30 minutes.
 All annotators will be compensated for their time with a \$20 gift card upon completion of the task.
+A.5.1 Study 1: verify the quality of generated images
 Task 1 Instructions
@@
 The difference between stereotypical clue and regular clue is that stereotypical would be something like baguettes for France or “naan” for India or specific clothing styles for some country whereas a regular clue is something that you are using to make your guess but you don’t know enough about your guess to know what stereotypical clues might be associated with it, for example, sand for island countries.
+A.5.2 Study 2: human performance on our cultural awareness benchmark
 Figure 19: Annotation Interface for Study 2
@@
 Then select at least one clue corresponding to your rating, similar to the first study.
-C.4 Additional Results
-$\Delta_{1}$ $\Delta_{2}$
-Greece $\rightarrow$ China $-8.48$ $5.07$
-Greece $\rightarrow$ India $-3.47$ $0.11$
-Greece $\rightarrow$ USA $-6.13$ $5.81$
-Table 3: Delta values for cultural adaptation from source country to target country. A successful cultural adaptation is indicated by $\Delta_{1}<0$ and $\Delta_{2}>0$.
 Performance of Human Study participants on the Cultural Awareness Task
@@
 Continent is the most generic label, and has very high accuracies from all participants.
+A.5.3 Study 3, 4: artifact hallucination and cultural relevance
+We provide an interface (Figure 20) that includes a multi-choice correct question answering setting, where we first ask a question about the artifacts that are present in the image and then ask if those artifacts are culturally relevant.
+For selecting these options, we sample randomly from all artifacts extracted for the accompanying image.
+For this study, we also provide the name of the concept class that the image is supposed to represent and the country, so that annotators have an idea about the object categories that they might be looking for.
+To account for cases where annotators might not be sure about any object and still choose one, in our analysis we only consider those cases where atleast two artifacts are marked as present.
+Responses to the second question is somewhat subjective as it depends on the cultural backgrounds of the annotators, but as we see in our analysis, the responses mostly agree and find more than half of the artifacts to be relevant.
+Figure 20: Annotation Interface for Study 3,4
+A.5.4 Study 5: people count associations
+We used an identical interface to the one for Study 3,4 to carry out this study, but instead of providing multiple correct options, we provide a direct True/False setting for annotators to answer if the count of people shown in the given image is approximately correct, i.e.
+falls in the correct count bucket (1-5, 5-10, more than 10).
+Our analysis reveals that people mostly agree with the count buckets generated, but on corroborating patterns with actual population statistics, the ordering of countries look slightly different, so we do not make any correlations about that in our analysis.
+A.5.5 Study 6, 7: layout preservation and cultural relevance after editing
+We use an interface (Figure 21) that shows the name of the source and target countries, the source image, and two edited images - one from our method CultureAdapt and one from cap-edit.
+We ask a somewhat subjective question regarding which image the annotator prefers, but responses vary widely for this question and do not correlate well across annotators, so we do not use these results for further analysis.
+For each of the edited images, we ask three questions - (1) if the edit maintains structural layout, (2) if the edit makes the image more culturally relevant to the target country and (3) if the edited image might be toxic for people from target country.
+All of these questions require a response on a Likert scale, which we design using a system of stars, where users can select anywhere between 1 and 5 stars.
+Less than 1% of images are marked toxic across all annotations and annotators agree strongly on this metric.
+We report the analysis for the other two metrics in the main paper.
+Figure 21: Annotation Interface for Study 6,7
+A.6 Additional Results
+A.6.1 Task 1 - Cultural awareness
+Income distribution performance for LLaVA
+We show the performance of LLaVA on the income quartile distribution basis for Dollar Street.
+Figure 22: We normalize income data from Dollar Street into region specific quartiles and plot corresponding accuracies for LLaVA.
+Confusion matrix set 1
+We show confusion matrices for both LLaVA and GPT-4V on Dalle Street images.
+Figure 23: Confusion matrices for GPT-4V and LLaVA on the cultural awareness task for Dalle Street images.
+Confusion matrix set 2
+We show confusion matrices for both LLaVA and GPT-4V on Dollar Street images.
+Figure 24: Confusion matrices for GPT-4V and LLaVA on the cultural awareness task for Dollar Street images.
+Confusion matrix set 3
+We show confusion matrices for both LLaVA and GPT-4V on MaRVL images.
+Figure 25: Confusion matrices for GPT-4V and LLaVA on the cultural awareness task for MaRVL images.
+A.6.2 Task 2 - Artifacts
 How many artifacts did we identify using GPT-4V
@@
 Further human filtering can be done to remove common words like table or mailbox or dresses to find unique and interesting associations like pretzels in Austria, zinnias in Bolivia and many more, some of which we report in Table LABEL:tab:artifact-interesting.
-Table 5: Artifact Statistics
+Figure 26: We identify more than 18,000 unique cultural artifacts across all countries as part of our second task, and then filter them to find salient ones.
+This figure shows strongest correlated artifacts for 20 randomly picked countries.
+Note that these associations are extracted from LMM generations and may not always be accurate.
 Austria Bangladesh Bolivia Brazil Bulgaria Burkina Faso Burundi Cambodia Cameroon Canada
@@
 no_adj 133 132 148 181 162 163 166 10035
-More examples of edits using CultureAdapt
-We include more examples of edits across different concept classes and source-target pairs using our CultureAdapt pipeline in Figure 14.
-As can be seen, the pipeline is only constrained by the two bottlenecks of object detection and diffusion based inpainting, which sometimes may detect objects incorrectly or not generate consistent images of human faces for example.
-Figure 12: We normalize income data from Dollar Street into region specific quartiles and plot corresponding accuracies for LLaVA.
-Figure 13: We identify more than 18,000 unique cultural artifacts across all countries as part of our second task, and then filter them to find salient ones.
-This figure shows strongest correlated artifacts for 20 randomly picked countries.
-Note that these associations are extracted from LMM generations and may not always be accurate.
-Figure 14: We show examples of edits made using our CultureAdapt pipeline across 4 different concept classes and 12 pairs of unique source, target combinations to illustrate both cases where our pipeline excels and also where it is limited by the parts it is composed of.
-For all of these edits, our metric success criteria of $\Delta_{1}<0$ and $\Delta_{2}>0$ is satisfied.
-Figure 15: Confusion matrices for GPT-4V and LLaVA on the cultural awareness task for Dalle Street images.
-Figure 16: Confusion matrices for GPT-4V and LLaVA on the cultural awareness task for Dollar Street images.
-Figure 17: Confusion matrices for GPT-4V and LLaVA on the cultural awareness task for MaRVL images.
+Table 4: Salient Artifact Statistics (adj indicates descriptors like color, etc are part of the artifact name, whereas no_adj indicates the artifact name does not have such descriptors)
+Color associations on DALL-E 3 images
+We find that countries are more likely to be associated with particular colors, with some showing prominently strong associations.
 Figure 27: We explore how countries are distributed on a color spectrum by first calculating a global average RGB vector for Dalle Street images and then defining deltas along each axes aggregated at the country level.
 Takeaway: We find interesting associations - Greece is strongly correlated with blue, Burkina Faso with red.
+People-count associations on DALL-E 3 images
+Distributions validated by humans do not always correlate with actual population statistics.
 Figure 28: Here, we look at buckets of people counts in Dalle Street images aggregated at the country level, each of the subplots representing one bucket.
 Takeaway: Counts of people in images may not always accurately reflect population densities of the corresponding countries to scale.
-(a) The counts of answers obtained from our human study about the appropriateness of generated images shows that mostly people agree or are neutral, with only 2 disagreements overall.
-(b) We measure accuracy for the cultural awareness task at the country, subregion and region level and find increasing performance in this order across our annotator pool.
-Figure 20: (a) Appropriateness of generated images as perceived by human study participants.
-(b) Accuracy of cultural awareness at various geographical levels across our annotator pool for a subset of randomly sampled Dalle Street images.
+Interesting associations
+We show examples of interesting associations identified by models and humans at the country level, for our artifact extraction task.
 Table 5: Interesting associations and their explanations for various countries. Note that these associations are extracted from LMM generations and may not always be accurate.
@@
 Zimbabwe thatched huts, african clothing, animal carvings
-C.5 LLM hyperparameters
-We discuss the generation settings we used for our experiments, and also the associated costs and hardware.
-Generation settings
-  • DALL-E 3 images were generated for vivid and natural settings for standard quality and size $1024$ x $1024$
-  • GPT-4 and GPT-4V generations were obtained for temperature $=0.7$, top_p $=0.95$, no frequency or presence penalty, no stopping condition other than the maximum number of tokens to generate, max_tokens $=300$.
-  • LLaVA generations were obtained for temperature $=1.0$ and top_p $=1.0$, no penalties, and max_tokens $=128$.
-The reason for using a slightly higher temperature and top_p is to have more consistent outputs.
-In our initial experiments, LLaVA did not perform as well in terms of following instructions at the same temperatue setting as GPT-4V.
-  • For Grounding DINO, we use ShilongLiu/GroundingDINO from Hugging Face and set both box and text thresholds to 0.25 for the grounded box generations.
-  • For Stable Diffusion, we use stabilityai/stable-diffusion-2-inpainting from Hugging Face, and replace the autoencoder with stabilityai/sd-vae-ft-mse.
-We also use a DPMSolverMultistepScheduler for speeding up the generation process.
-We add "intricate details.
-4k.
-high resolution.
-high quality." to the end of our prompt to get high quality images.
-Computation budget
-  • We spent about \$$\$800$ in total for DALL-E 3 generations. This was funded by a grant from Microsoft Azure.
-  • We spent about \$$\$700$ in total for GPT-4V vision-preview and GPT-4 turbo inference and across all experiments.
-  • For experiments with LLaVA, Stable Diffusion and Grounding DINO, we used a single instance of a Multi-Instance A100 GPU with 40GB of GPU memory, 3/7 fraction of Streaming Multiprocessors, 2 NVIDIA Decoder hardware units, 4/8 L2 cache size, and 1 node.
-  • Total emissions for API based models are estimated to be 25 kgCO2 eq, of which 100 percents were directly offset by the cloud provider.
-Total emissions for our on-premise GPU usage is estimated to be less than 5 kgCO2 eq.
-Estimations were conducted using the MachineLearning Impact calculator Lacoste et al.
-(2019).
+A.6.3 Task 3 - Edits
+More examples of edits using CultureAdapt
+We include more examples of edits across different concept classes and source-target pairs using our CultureAdapt pipeline in Figure 29.
+As can be seen, the pipeline is only constrained by the two bottlenecks of object detection and diffusion based inpainting, which sometimes may detect objects incorrectly or not generate consistent images of human faces for example.
+Figure 29: We show examples of edits made using our CultureAdapt pipeline across 4 different concept classes and 12 pairs of unique source, target combinations to illustrate both cases where our pipeline excels and also where it is limited by the parts it is composed of.
+For all of these edits, our metric success criteria of $\Delta_{1}<0$ and $\Delta_{2}>0$ is satisfied.
+CultureAdapt quantitative metrics
+Source-Target Pair Similarity SSIM Metric $M_{1}$ Metric $M_{2}$
+cap-edit CultureAdapt cap-edit CultureAdapt cap-edit CultureAdapt
+Brazil-India 0.96 0.93 0.72 0.68 0.95 0.95
+Brazil-Nigeria 0.96 0.93 0.62 0.47 0.89 0.91
+Brazil-Turkey 0.96 0.93 0.69 0.52 0.95 0.94
+Brazil-USA 0.96 0.93 0.28 0.35 0.91 0.83
+Average 0.96 0.93 0.58 0.51 0.93 0.91
+India-Brazil 0.97 0.94 0.63 0.60 0.96 0.90
+India-Nigeria 0.96 0.94 0.69 0.68 0.94 0.90
+India-Turkey 0.97 0.94 0.56 0.57 0.92 0.89
+India-USA 0.96 0.94 0.67 0.63 0.93 0.88
+Average 0.97 0.94 0.63 0.62 0.94 0.89
+Nigeria-Brazil 0.96 0.92 0.29 0.41 0.76 0.85
+Nigeria-India 0.96 0.92 0.62 0.67 0.87 0.93
+Nigeria-Turkey 0.96 0.91 0.66 0.63 0.91 0.93
+Nigeria-USA 0.96 0.92 0.50 0.60 0.90 0.91
+Average 0.96 0.92 0.51 0.58 0.86 0.90
+Turkey-Brazil 0.97 0.94 0.46 0.41 0.89 0.84
+Turkey-India 0.97 0.95 0.62 0.59 0.88 0.88
+Turkey-Nigeria 0.97 0.94 0.67 0.64 0.93 0.90
+Turkey-USA 0.96 0.94 0.29 0.43 0.88 0.89
+Average 0.97 0.94 0.51 0.51 0.89 0.88
+USA-Brazil 0.97 0.94 0.53 0.46 0.88 0.89
+USA-India 0.98 0.94 0.18 0.62 0.58 0.91
+USA-Nigeria 0.98 0.94 0.20 0.46 0.61 0.84
+USA-Turkey 0.98 0.94 0.21 0.46 0.64 0.87
+Average 0.97 0.94 0.28 0.50 0.68 0.88
+Overall Average 0.97 0.94 0.50 0.54 0.85 0.89
+Table 7: Mean Similarity (SSIM) Scores, Metric $M_{1}$, and Metric $M_{2}$ for cap-edit and CultureAdapt grouped by Source and Target Country